### v1

# Double Yangians and quantum vertex algebras, I

Abstract.
For any simply-laced generalized Cartan matrix $A$, we introduce an algebra $\widehat{\mathcal{DY}}(A)$, which is essentially the centrally extended double Yangian when $A$ is of finite type, and we give a new field (current) presentation of $\widehat{\mathcal{DY}}(A)$. Among the main results, for any $\ell\in{\mathbb{C}}$ we construct a universal vacuum $\widehat{\mathcal{DY}}(A)$-module $\mathcal{V}_{A}(\ell)$ of level $\ell$, prove that there exists a natural $\hbar$-adic weak quantum vertex algebra structure on $\mathcal{V}_{A}(\ell)$, and give an isomorphism between the category of restricted $\widehat{\mathcal{DY}}(A)$-modules of level $\ell$ and the category of $\mathcal{V}_{A}(\ell)$-modules.
Key words and phrases:
Quantum vertex algebra, affine vertex algebra, centrally extended double Yangian
2010 Mathematics Subject Classification:
17B69
1Partially supported by NSF of China (No. 12371027, No. 12471029).
1. Introduction
It is well known that vertex algebras are naturally associated to affine Lie algebras, giving an important class of examples. In the general vertex algebra theory, an important subject arguably is establishing and exploring natural associations of quantum vertex algebras (in some sense) to quantum affine algebras and double Yangians.
As a fundamental work, Etingof and Kazhdan (see [EK]) introduced a theory of quantum vertex operator algebras and constructed a family of examples as formal deformations of the (universal) affine vertex algebras of type $A$. Since then, this theory has been extensively studied and developed in various directions.
In [Li2, Li3], notions of nonlocal vertex algebra and (weak) quantum vertex algebra were introduced and systematically studied, where nonlocal vertex algebras are vertex algebra analogues of noncommutative associative algebras while (weak) quantum vertex algebras are generalizations of vertex superalgebras (including vertex algebras). As slight generalizations of quantum vertex operator algebras in the sense of Etingof-Kazhdan, $\hbar$-adic (weak) quantum vertex algebras, which are formal deformations of (weak) quantum vertex algebras, were studied in [Li5]. In particular, a conceptual construction of $\hbar$-adic weak quantum vertex algebras and their modules was established by using $\mathcal{S}$}-local” sets of fields (vertex operators). As an application, $\hbar$-adic quantum vertex algebras were associated to the centrally extended double Yangian of $\mathfrak{sl}_{2}$.
As for quantum affine algebras, a theory of what were called $\phi$-coordinated (quasi) modules for (weak) quantum vertex algebras was developed in [Li6] and a conceptual association of quantum vertex algebras to quantum affine algebras in terms of $\phi$-coordinated (quasi) modules was obtained. (Quantum vertex algebras were explicitly associated to some other algebras, e.g., to the deformed Virasoro algebra $\mathcal{V}_{p,q}$ with q=1q=-1 in [Li7] and to the Ding-Iohara algebra with $c=0$ in [LTW].)
On the other hand, Butorac, Jing and Kožić (see [BJK]) extended Etingof-Kazhdan’s construction to types $B$, $C$ and $D$ (with rational $R$-matrices), by using the $R$-matrix type realization of double Yangians in [JYL]. It was proved in [Koz2] that modules for these quantum vertex operator algebras are in one-to-one correspondence with restricted modules for the corresponding double Yangians. Recently, by utilizing the $R$-matrix presentation of quantum affine algebras (see [DF, JLM1, JLM2]), Kožić (see [Koz1, Koz2]) constructed the quantum vertex operator algebras associated with trigonometric $R$-matrices of types $A$, $B$, $C$ and $D$, and established a one-to-one correspondence between $\phi$-coordinated modules and restricted modules for quantum affine algebras.
In a previous paper [JKLT2] joint with Jing and Tan, by using the smash product construction of nonlocal vertex algebras (see [Li4]), we constructed a family of $\hbar$-adic quantum vertex algebras by explicitly deforming lattice vertex algebras. In [Kon], mainly by deforming affine vertex algebras the first named author of this paper constructed a family of $\hbar$-adic quantum affine vertex algebras and their $\phi$-coordinated quasi modules.
In this current paper, we continue studying centrally extended double Yangians in the context of $\hbar$-adic quantum vertex algebras and their modules, using the generator-and-defining-relation approach. We first introduce an algebra $\widehat{\mathcal{DY}}(A)$ for any simply-laced generalized Cartan matrix $A$, which essentially coincides with the centrally extended double Yangian when $A$ is of finite type, and we then give a new field (current) presentation of $\widehat{\mathcal{DY}}(A)$. Among the main results, for any $\ell\in{\mathbb{C}}$, we construct a universal vacuum $\widehat{\mathcal{DY}}(A)$-module $\mathcal{V}_{A}(\ell)$ of level $\ell$, prove that there exists a natural $\hbar$-adic weak quantum vertex algebra structure on $\mathcal{V}_{A}(\ell)$, and give an isomorphism between the category of restricted $\widehat{\mathcal{DY}}(A)$-modules of level $\ell$ and the category of $\mathcal{V}_{A}(\ell)$-modules.
In the following, we discuss some technical details and mention some relevant results. Note that for any (possibly infinite-dimensional) Lie algebra $\mathfrak{g}$} equipped with a non-degenerate invariant bilinear form, a general affine Lie algebra $\widehat{{\mathfrak{g}}}$ can be defined explicitly in the sense that $\widehat{{\mathfrak{g}}}$ is linearly spanned by the standard central element $\bf k$} and the coefficients of the generating functions $a(z)$ for $a\in{\mathfrak{g}}$, and commutators $[a(x),b(z)]$ are “linearly closed.” Unlike affine Lie algebras, centrally extended double Yangians (and as well as quantum affine algebras) are defined by Chevalley generators and relations, including Serre relations for the high rank case. As for $\mathfrak{sl}_{2}$, the centrally extended double Yangian is defined in terms of four generating functions $E(z),F(z),H^{+}(z),H^{-}(z)$. In [Li5], these original currents were used as fields to generate an $\hbar$-adic quantum vertex algebra, especially, with $H^{+}(z)$ (annihilation) and $H^{-}(z)$ (creation) considered as two fields. Consequently, the resulted $\hbar$-adic quantum vertex algebra is a formal deformation of the affine vertex algebra associated to a $4$-dimensional Lie algebra. Naturally, one hopes to have an $\hbar$-adic quantum vertex algebra which is a formal deformation of the affine vertex algebra associated to $\mathfrak{sl}_{2}$.
With this as one of the motivations, for the algebra $\widehat{\mathcal{DY}}(A)$ we introduce a new set of generating fields denoted by $x_{i,\mathcal{Y}}^{\pm}(z),h_{i,\mathcal{Y}}(z)$ for $i\in I$, where in particular the modified Cartan annihilation and creation currents are combined. Then we give a characterization of $\widehat{\mathcal{DY}}(A)$ in terms of this new set of fields. In the new presentation, the same vertex-operator Serre relations are proved to hold, and $\mathcal{S}$}-commutators” $x_{i,\mathcal{Y}}^{+}(z_{1}),x_{i,\mathcal{Y}}^{-}(z_{2})]_{{\mathcal{S}}}$ replace the usual Lie bracket. For proving the equivalence, we give two characterizations of the vertex-operator Serre relation. On the other hand, to deal with the $\mathcal{W}$-algebra type nonlinear relations in the $\mathcal{S}$}-commutators, we prove and use a vertex-operator iterate formula. Just as with the association of vertex algebras to affine Lie algebras, we need to construct “universal vacuum modules” for $\widehat{\mathcal{DY}}(A)$ and show they have a natural $\hbar$-adic quantum vertex algebra structure. While the construction of universal vacuum modules $\mathcal{V}_{A}(\ell)$ is mostly tautological, to get topologically free $\mathbb{C}$$\hbar$]]-modules in several stages, we make use of a new gadget—notions of strong submodule and ideal.
In a sequel, we plan to determine the structure of the universal vacuum $\widehat{\mathcal{DY}}(A)$-module $\mathcal{V}_{A}(\ell)$ and hopefully show that ${\mathcal{V}}_{A}(\ell)/\hbar{\mathcal{V}}_{A}(\ell)$ is a universal vacuum $\widehat{{\mathfrak{g}}}-$module of level $\ell$ with ${\mathfrak{g}}={\mathfrak{g}}(A)$ for $A$ of a finite type. This implies that the $\hbar$-adic weak quantum vertex algebra ${\mathcal{V}}_{A}(\ell)$ is a non-degenerate quantum vertex operator algebra in the sense of Etingof and Kazhdan.
This paper is organized as follows. Section 2 is preliminary, in which we recall some basics on nonlocal vertex algebras and (weak) quantum vertex algebras. In Section 3, we recall some basic results on $\hbar$-adic (weak) quantum vertex algebras. In Section 4, we establish several technical results including a vertex operator iterate formula and equivalent characterizations of the vertex-operator Serre relation. In Section 5, we study algebra $\widehat{\mathcal{DY}}(A)$ and $\hbar$-adic weak quantum vertex algebra ${\mathcal{V}}_{A}(\ell)$, and present the natural connection of $\widehat{\mathcal{DY}}(A)$-modules with ${\mathcal{V}}_{A}(\ell)$-modules.
In this paper, with $\mathbb{C}$} (complex numbers) as our scalar field we use the formal variable notations and conventions as established in [FLM] and [FHL]. In particular, ${\mathbb{C}}(x)$ and ${\mathbb{C}}(x,y)$ denote the fields of rational functions. We also use $\partial_{x}$ for the formal differential operator $\frac{\partial}{\partial x}$. In addition, we use $\mathbb{Z}$$_{+}$ and $\mathbb{N}$} for the sets of positive integers and nonnegative integers, respectively.
2. Vertex algebras and quantum vertex algebras
In this section, we review some basic notions and results, to introduce some frequently used terminologies and notations.
We begin by recalling the definition of a vertex algebra. A vertex algebra is a vector space $V$ equipped with a linear map
$Y(\cdot,x):\$$V\to(\textnormal{{End}}V)[[x,x^{-1}]]$
(2.1) $v\mapsto Y(v,x)=\sum_{n\in{\mathbb{Z}}}v_{n}x^{-n-1}\ \ (\textnormal{{where }}v_{n}\in\textnormal{{End}}V)$
and equipped with a vector ${{\bf 1}}\in V$, called the vacuum vector, such that
(2.2) $Y(u,x)v\in V((x))\quad\text{for }u,v\in V,$
(2.3) $Y({{\bf 1}},x)=1\ \ \ (\text{the identity operator on }V),$
(2.4) $Y(v,x){{\bf 1}}\in V[[x]]\quad\textnormal{{and}}\quad\lim_{x\to 0}Y(v,x){{\bf 1}}=v\quad\textnormal{{for }}v\in V,$
and such that the following Jacobi identity holds for $u,v,w\in V$:
$x_{0}^{-1}\delta\!\left(\frac{x_{1}-x_{2}}{x_{0}}\right)$ $Y(u,x_{1})Y(v,x_{2})w-x_{0}^{-1}\delta\!\left(\frac{x_{2}-x_{1}}{-x_{0}}\right)\!Y(v,x_{2})Y(u,x_{1})w$
(2.5) $=x_{1}^{-1}\delta\!\left(\frac{x_{2}+x_{0}}{x_{1}}\right)\!Y(Y(u,x_{0})v,x_{2})w.$
A consequence of the Jacobi identity is Borcherds’ commutator formula
(2.6) $[Y(u,x_{1}),Y(v,x_{2})]=\sum_{n\geq 0}Y(u_{n}v,x_{2})\frac{1}{n!}\left(\frac{\partial}{\partial x_{2}}\right)^{n}x_{1}^{-1}\delta\!\left(\frac{x_{2}}{x_{1}}\right)\!.$
This formula further implies the weak commutativity, namely locality: For any $u,v\in V$, there exists $k\in{\mathbb{N}}$ such that
(2.7) $(x_{1}-x_{2})^{k}Y(u,x_{1})Y(v,x_{2})=(x_{1}-x_{2})^{k}Y(v,x_{2})Y(u,x_{1}).$
The following analogue of the notion of (noncommutative) associative algebra (see [BK], [Li3]) plays a key role in the study of quantum vertex algebras:
Definition 2.1.
A nonlocal vertex algebra is defined by using all the axioms that define a vertex algebra except the Jacobi identity axiom which is replaced by the weak associativity: For any $u,v,w\in V$, there exists $l\in{\mathbb{N}}$ such that
(2.8) $(x_{0}+x_{2})^{l}Y(u,x_{0}+x_{2})Y(v,x_{2})w=(x_{0}+x_{2})^{l}Y(Y(u,x_{0})v,x_{2})w.$
Let $V$ be a nonlocal vertex algebra. Denote by $\mathcal{D}$ the linear operator on $V$ given by
(2.9) $\mathcal{D}(v)=v_{-2}{\bf 1}=\left(\frac{d}{dz}Y(v,z){\bf 1}\right)|_{z=0}\quad\text{ for }v\in V.$
The following are the basic properties:
(2.10) $[\mathcal{D},Y(v,x)]=Y(\mathcal{D}(v),x)=\frac{d}{dx}Y(v,x),$
(2.11) $e^{z\mathcal{D}}Y(v,x)e^{-z\mathcal{D}}=Y(e^{z\mathcal{D}}v,x)=Y(v,x+z),$
(2.12) $Y(v,x){\bf 1}=e^{x\mathcal{D}}v\quad\quad\text{ for }v\in V.$
An important class of nonlocal vertex algebras consists of what are called weak quantum vertex algebras (see [Li2]).
Definition 2.2.
A weak quantum vertex algebra is defined by using all the axioms that define a vertex algebra $V$ except the Jacobi identity which is replaced with the property that for any $u,v\in V$, there exist (finitely many)
$u^{(i)},v^{(i)}\in V,\ f_{i}(x)\in{\mathbb{C}}((x))\ \ \text{ for }1\leq i\leq r$
such that the following $\mathcal{S}$}-Jacobi identity holds:
(2.13) $x_{0}^{-1}\delta\!\left(\frac{x_{1}-x_{2}}{x_{0}}\right)\!Y(u,x_{1})Y(v,x_{2})$
$\quad-x_{0}^{-1}\delta\!\left(\frac{x_{2}-x_{1}}{-x_{0}}\right)\sum_{i=1}^{r}f_{i}(x_{2}-x_{1})Y(v^{(i)},x_{2})Y(u^{(i)},x_{1})$
[…]
The $\mathcal{S}$}-Jacobi identity axiom above is equivalent to the combination of the weak associativity and the $\mathcal{S}$}-locality (see [EK]): For any $u,v\in V$, there exist
$u^{(i)},v^{(i)}\in V,\ f_{i}(x)\in{\mathbb{C}}((x))\ \ \text{ for }1\leq i\leq r$
and a nonnegative integer $k$ such that
(2.14) $(x_{1}-x_{2})^{k}Y(u,x_{1})Y(v,x_{2})=(x_{1}-x_{2})^{k}\sum_{i=1}^{r}f_{i}(x_{2}-x_{1})Y(v^{(i)},x_{2})Y(u^{(i)},x_{1}).$
Remark 2.3.
Let $V$ be a weak quantum vertex algebra. From definition, there is a linear map ${\mathcal{S}}(x):\ V\otimes V\rightarrow V\otimes V\otimes{\mathbb{C}}((x))$ such that for $u,v\in V$, (2.13) holds with
(2.15) ${\mathcal{S}}(x)(v\otimes u)=\sum_{i=1}^{r}v^{(i)}\otimes u^{(i)}\otimes f_{i}(x)\in V\otimes V\otimes{\mathbb{C}}((x)).$
Such a linear map ${\mathcal{S}}(x)$ is called an $\mathcal{S}$}-locality operator of $V$.
Let $V$ be a general nonlocal vertex algebra. Following [EK], let
(2.16) $Y(x):\ V\otimes V\rightarrow V((x))$
denote the canonical linear map associated to the vertex-operator map $Y(\cdot,x)$ with
$Y(x)(u\otimes v)=Y(u,x)v\ \ \textnormal{{ for }}u,v\in V.$
A rational quantum Yang-Baxter operator on a vector space $U$ is a linear map
${\mathcal{S}}(x):\ U\otimes U\rightarrow U\otimes U\otimes{\mathbb{C}}((x)),$
satisfying the following rational quantum Yang-Baxter equation:
(2.17) ${\mathcal{S}}^{12}(z){\mathcal{S}}^{13}(z+w){\mathcal{S}}^{23}(w)={\mathcal{S}}^{23}(w){\mathcal{S}}^{13}(z+w){\mathcal{S}}^{12}(z)$
on $U\otimes U\otimes U$. It is said to be unitary if
(2.18) ${\mathcal{S}}^{21}(x){\mathcal{S}}(-x)=1,$
where ${\mathcal{S}}^{21}(x)=\sigma{\mathcal{S}}(x)\sigma$ with $\sigma$ denoting the flip operator on $U\otimes U$.
Definition 2.4.
A quantum vertex algebra is a weak quantum vertex algebra $V$ equipped with a unitary rational quantum Yang-Baxter operator ${\mathcal{S}}(x)$ on $V$, such that ${\mathcal{S}}(x)$ is an $\mathcal{S}$}-locality operator of $V$ and such that
(2.19) $\hskip 42.67912pt{\mathcal{S}}(x)(v\otimes{\bf 1})=v\otimes{\bf 1}\quad\text{ for }v\in V,$
(2.20) $[\mathcal{D}\otimes 1,{\mathcal{S}}(x)]=-\frac{d}{dx}{\mathcal{S}}(x)\quad\text{(the \emph{shift condition})},$
(2.21) ${\mathcal{S}}(z)(Y(w)\otimes 1)=(Y(w)\otimes 1){\mathcal{S}}^{23}(z){\mathcal{S}}^{13}(z+w)\quad\text{ (the \emph{hexagon identity})}$
on $V\otimes V\otimes V$.
The following notion is due to Etingof and Kazhdan (see [EK]):
Definition 2.5.
A nonlocal vertex algebra $V$ is said to be non-degenerate if for every positive integer $n$, the linear map $Z_{n}:\ V^{\otimes n}\otimes{\mathbb{C}}((x_{1}))\cdots((x_{n}))\rightarrow V((x_{1}))\cdots((x_{n}))$ is injective, where
(2.22) $Z_{n}(v^{(1)}\otimes\cdots\otimes v^{(n)}\otimes f)=f(x_{1},\dots,x_{n})Y(v^{(1)},x_{1})\cdots Y(v^{(n)},x_{n}){\bf 1}.$
The following result, which was formulated in [Li2], is essentially due to [EK]:
Proposition 2.6.
Let $V$ be a weak quantum vertex algebra. Assume that $V$ is non-degenerate. Then there exists a unique $\mathcal{S}$}-locality operator ${\mathcal{S}}(x)$. Furthermore, ${\mathcal{S}}(x)$ is a unitary rational quantum Yang-Baxter operator and $(V,{\mathcal{S}}(x))$ is a quantum vertex algebra.
Remark 2.7.
Note that an $\mathcal{S}$}-locality operator is part of a quantum vertex algebra structure by definition. In view of Proposition 2.6, the term “non-degenerate quantum vertex algebra” (without specifying an $\mathcal{S}$}-locality operator) is unambiguous.
Definition 2.8.
Let $V$ be a nonlocal vertex algebra. A $V$-module is a vector space $W$ equipped with a linear map
$Y_{W}(\cdot,x):\$$V\to(\textnormal{{End}}W)[[x,x^{-1}]];\quad v\mapsto Y_{W}(v,x),$
satisfying the conditions that $Y_{W}({\bf 1},x)=1_{W}$ (the identity operator on $W$),
$Y_{W}(v,x)w\in W((x))\quad\text{ for }v\in V,\ w\in W,$
and that for $u,v\in V,\ w\in W$, there exists $l\in{\mathbb{N}}$ such that
(2.23) $(x_{0}+x_{2})^{l}Y_{W}(u,x_{0}+x_{2})Y_{W}(v,x_{2})w=(x_{0}+x_{2})^{l}Y_{W}(Y(u,x_{0})v,x_{2})w.$
Let $W$ be a general vector space (over $\mathbb{C}$}). Set
(2.24) ${\mathcal{E}}(W)=\operatorname{Hom}(W,W((x)))\subset(\operatorname{End}W)[[x,x^{-1}]].$
A typical element of ${\mathcal{E}}(W)$ is the identity operator on $W$, denoted by $1_{W}$.
Definition 2.9.
A finite sequence $a^{[1]}(x),\dots,a^{[r]}(x)$ in ${\mathcal{E}}(W)$ is said to be compatible if there exists a nonnegative integer $k$ such that
(2.25) $\left(\prod_{1\leq i<j\leq r}(x_{i}-x_{j})^{k}\right)a^{[1]}(x_{1})\cdots a^{[r]}(x_{r})\in\operatorname{Hom}(W,W((x_{1},\dots,x_{r}))).$
A subset $U$ of ${\mathcal{E}}(W)$ is said to be compatible if every finite sequence in $U$ is compatible.
Let $(a(x),b(x))$ be a compatible ordered pair in ${\mathcal{E}}(W)$. Define $a(x)_{n}b(x)\in{\mathcal{E}}(W)$ for $n\in{\mathbb{Z}}$ in terms of generating function
$Y_{{\mathcal{E}}}(a(x),z)b(x)=\sum_{n\in{\mathbb{Z}}}a(x)_{n}b(x)z^{-n-1}$
by
(2.26) $Y_{{\mathcal{E}}}(a(x),z)b(x)=z^{-k}\left((x_{1}-x)^{k}a(x_{1})b(x)\right)|_{x_{1}=x+z},$
where $k$ is any nonnegative integer such that
$(x_{1}-x)^{k}a(x_{1})b(x)\in\operatorname{Hom}(W,W((x,x_{1}))).$
A compatible subspace $U$ of ${\mathcal{E}}(W)$ is said to be $Y_{\mathcal{E}}$-closed if
[…]
The following result was obtained in [Li2]:
Theorem 2.10.
Let $V$ be a $Y_{{\mathcal{E}}}$-closed compatible subspace of ${\mathcal{E}}(W)$ with $1_{W}\in V$. Then $(V,Y_{\mathcal{E}},1_{W})$ is a nonlocal vertex algebra and $W$ is a $V$-module with $Y_{W}(a(x),z)=a(z)$ for $a(x)\in V$. On the other hand, for any compatible subset $U$ of ${\mathcal{E}}(W)$, there exists a (unique) smallest $Y_{\mathcal{E}}$-closed compatible subspace ${\langle}U{\rangle}$ such that $U\cup\{1_{W}\}\subset{\langle}U{\rangle}$, where
(2.27) ${\langle}U{\rangle}={\rm span}\{u^{(1)}(x)_{m_{1}}\cdots u^{(r)}(x)_{m_{r}}1_{W}\ |\ r\geq 0,\ u^{(i)}(x)\in U,\ m_{i}\in{\mathbb{Z}}\}.$
A subset $U$ of ${\mathcal{E}}(W)$ is said to be $\mathcal{S}$}-local if for any $a(x),b(x)\in U$, there exist
$c_{i}(x),\ d_{i}(x)\in U,\ f_{i}(x)\in{\mathbb{C}}((x))\ \textnormal{{ for }}1\leq i\leq r$
and a nonnegative integer $k$ such that
(2.28) $(x_{1}-x_{2})^{k}a(x_{1})b(x_{2})=(x_{1}-x_{2})^{k}\sum_{i=1}^{r}f_{i}(x_{2}-x_{1})c_{i}(x_{2})d_{i}(x_{1}).$
With this notion, we have (see [Li2]):
Proposition 2.11.
Let $U$ be an $\mathcal{S}$}-local subset of ${\mathcal{E}}(W)$. Then $U$ is compatible and ${\langle}U{\rangle}$ is a weak quantum vertex algebra.
3. $\hbar$-adic quantum vertex algebras and their modules
In this section, we recall some basic notions and results on $\hbar$-adic nonlocal vertex algebras and $\hbar$-adic (weak) quantum vertex algebras.
3.1. Topologically free $\mathbb{C}$$\hbar$]]-modules
Let $\hbar$ be a formal variable. A $\mathbb{C}$$\hbar$]]-module $V$ is said to be torsion-free if $\hbar v\neq 0$ for any nonzero vector $v\in V$, and said to be separated if $\cap_{n\geq 1}\hbar^{n}V=0$. For a $\mathbb{C}$$\hbar$]]-module $V$, using subsets $v+\hbar^{n}V$ for $v\in V$, $n\geq 1$ as the basis of open sets one obtains a topology on $V$, which is called the $\hbar$-adic topology. A $\mathbb{C}$$\hbar$]]-module $V$ is said to be $\hbar$-adically complete if every Cauchy sequence in $V$ with respect to the $\hbar$-adic topology has a limit in $V$. A $\mathbb{C}$$\hbar$]]-module $V$ is topologically free if $V=V_{0}[[\hbar]]$ for some $\mathbb{C}$}-subspace $V_{0}$ of $V$.
Let $U$ and $V$ be topologically free $\mathbb{C}$$\hbar$]]-modules. The $\hbar$-adically completed tensor product $U\widehat{\otimes}V$ of $U$ and $V$ is defined to be the $\hbar$-adic completion of $U\otimes V$. If $U=U_{0}[[\hbar]]$ and $V=V_{0}[[\hbar]]$, then $U\widehat{\otimes}V=(U_{0}\otimes V_{0})[[\hbar]]$.
The following basic facts can be either found (cf. [Ka]) or proved straightforwardly:
Lemma 3.1.
(1) A $\mathbb{C}$$\hbar$]]-module is topologically free if and only if it is torsion-free, separated, and $\hbar$-adically complete. (2) If $W$ is a topologically free $\mathbb{C}$$\hbar$]]-module, then $W=U[[\hbar]]$ for any $\mathbb{C}$}-subspace $U$ of $W$ such that $W=U\oplus\hbar W$ over $\mathbb{C}$}. (3) If $U$ is a $\mathbb{C}$$\hbar$]]-submodule of a torsion-free and separated $\mathbb{C}$$\hbar$]]-module and if $U$ itself is $\hbar$-adically complete, then $U$ is topologically free.
Lemma 3.2.
Let $W$ be a topologically free $\mathbb{C}$$\hbar$]]-module and let $K$ be a $\mathbb{C}$}-subspace of $W$. Define $K[[\hbar]]^{\prime}$ to be the image of $K[[\hbar]]$ in $W$ under the canonical $\mathbb{C}$$\hbar$]]-module morphism from $K[[\hbar]]$ to $W$. Then $K[[\hbar]]^{\prime}$ is topologically free. On the other hand, a $\mathbb{C}$$\hbar$]]-submodule $K$ of $W$ itself is $\hbar$-adically complete if and only if $K=K[[\hbar]]^{\prime}$.
The following result can be found in [Ka]:
Lemma 3.3.
Let $W_{1},W_{2}$ be topologically free $\mathbb{C}$$\hbar$]]-modules and let $\psi:W_{1}\rightarrow W_{2}$ be a continuous $\mathbb{C}$$\hbar$]]-module map. If the derived $\mathbb{C}$}-linear map $\bar{\psi}:W_{1}/\hbar W_{1}\rightarrow W_{2}/\hbar W_{2}$ is one-to-one (resp. onto), then $\psi$ is one-to-one (resp. onto).
We also have the following simple facts (see [Li5]; the proof of Proposition 3.7):
Lemma 3.4.
Let $U$ be a $\mathbb{C}$$\hbar$]]-submodule of a $\mathbb{C}$$\hbar$]]-module $W$. Set
(3.1) $[U]=\{w\in W\ |\ \hbar^{n}w\in U\ \text{ for some }n\in{\mathbb{N}}\}.$
Then $\left[[U]\right]=[U]$. On the other hand, if $[U]=U$, then $U\cap\hbar^{n}W=\hbar^{n}U$ for $n\in{\mathbb{N}}$.
Lemma 3.5.
Let $U$ be a $\mathbb{C}$$\hbar$]]-submodule of a topologically free $\mathbb{C}$$\hbar$]]-module $W$ such that $[U]=U$. Then $\left[\bar{U}\right]=\bar{U}$, $\overline{\left[\bar{U}\right]}=\left[\bar{U}\right]$, and $\bar{U}$ is topologically free.
Lemma 3.6.
Let $U$ and $V$ be topologically free $\mathbb{C}$$\hbar$]]-modules and let $\psi:U\rightarrow V$ be a $\mathbb{C}$$\hbar$]]-module map. Then $\ker(\psi)]=\ker(\psi)$, $\overline{\ker(\psi)}=\ker(\psi)$, and $\ker(\psi)$ is topologically free.
Lemma 3.7.
Let $A$ be an associative algebra over $\mathbb{C}$$\hbar$]], which is topologically free as a $\mathbb{C}$$\hbar$]]-module.
1) For any left (right) ideal $J$ of $A$ as a $\mathbb{C}$$\hbar$]]-algebra, $[J]$ is a left (right) ideal.
2) For any left (right) ideal $J$ of $A$ such that $[J]=J$, $\overline{J}$ is a left (right) ideal of $A$ such that $\overline{J}]=\overline{J}$ and $\overline{J}$ is topologically free.
Lemma 3.8.
Let $W$ be a topologically free $\mathbb{C}$$\hbar$]]-module and let $U$ be a $\mathbb{C}$$\hbar$]]-submodule such that $[U]=U$ and $U$ is $\hbar$-adically complete. Then there exist a $\mathbb{C}$}-subspace $W_{0}$ of $W$ and a $\mathbb{C}$}-subspace $U_{0}$ of $W_{0}$ such that $W=W_{0}[[\hbar]]$ and $U=U_{0}[[\hbar]]$. Furthermore, $W/U=(W_{0}/U_{0})[[\hbar]]$. In particular, $W/U$ is topologically free.
Proof.
As a $\mathbb{C}$$\hbar$]]-submodule of $W$, $U$ is torsion-free and separated. In addition, $U$ is $\hbar$-adically complete by assumption. Then $U$ is topologically free, i.e., $U=U_{0}[[\hbar]]$ for some $\mathbb{C}$}-subspace $U_{0}$ of $U$. Noticing that $U\cap\hbar W=\hbar U$ by Lemma 3.4, we have $U_{0}\cap\hbar W=U_{0}\cap U\cap\hbar W=U_{0}\cap\hbar U=0$. Then there exists a $\mathbb{C}$}-subspace $W_{0}^{\prime}$ of $W$ such that $W=W_{0}^{\prime}\oplus U_{0}\oplus\hbar W$ over $\mathbb{C}$}. Set $W_{0}=W_{0}^{\prime}+U_{0}$. Then $U_{0}\subset W_{0}$ and $W=W_{0}[[\hbar]]$. Next, we prove $W/U=(W_{0}/U_{0})[[\hbar]]$. The natural $\mathbb{C}$}-linear map $\theta:W_{0}\rightarrow W_{0}/U_{0}$ gives a surjective $\mathbb{C}$$\hbar$]]-morphism $\tilde{\theta}:W=W_{0}[[\hbar]]\rightarrow(W_{0}/U_{0})[[\hbar]]$. It is clear that $U=U_{0}[[\hbar]]\subset\ker\tilde{\theta}$, so we have a surjective $\mathbb{C}$$\hbar$]]-morphism $\bar{\theta}:W/U\rightarrow(W_{0}/U_{0})[[\hbar]]$. Suppose
$v+U=(v_{0}+\hbar v_{1}+\hbar^{2}v_{2}+\cdots)+U\in\ker\bar{\theta}$
with $v_{n}\in W_{0}$ for $n\geq 0$. Then $\theta(v_{0})+\hbar\theta(v_{1})+\hbar^{2}\theta(v_{2})+\cdots=0$ in $(W_{0}/U_{0})[[\hbar]]$, which implies $\theta(v_{n})=0$, i.e., $v_{n}\in U_{0}$ for all $n\geq 0$. Thus $v\in U$. This proves that $\bar{\theta}$ is also injective, and hence $\bar{\theta}$ is a $\mathbb{C}$$\hbar$]]-isomorphism. ∎
Let $M$ be a $\mathbb{C}$$\hbar$]]-module. For a sequence $\{w(n)\}_{n\in{\mathbb{N}}}$ in $M$, we write $\lim_{n\rightarrow\infty}w(n)=0$ to signify that for every $r\in{\mathbb{Z}}_{+}$, there exists $k\in{\mathbb{Z}}_{+}$ such that $w(n)\in\hbar^{r}M$ for all $n\geq k$. Define
(3.2) $M_{\hbar}((x))=\left\{\left.\psi(x)=\sum_{n\in{\mathbb{Z}}}\psi(n)x^{n}\in M[[x,x^{-1}]]\ \right|\ \lim_{n\rightarrow\infty}\psi(-n)=0\right\}.$
If $M=M_{0}[[\hbar]]$ is a topologically free $\mathbb{C}$$\hbar$]]-module, then $M_{\hbar}((x))=(M_{0}((x)))[[\hbar]]$, which is topologically free.
Let $W=W_{0}[[\hbar]]$ be a topologically free $\mathbb{C}$$\hbar$]]-module. Then $\operatorname{End}W=(\operatorname{End}W_{0})[[\hbar]]$, which is a topologically free $\mathbb{C}$$\hbar$]]-module. Define
(3.3) ${\mathcal{E}}_{\hbar}(W)=\operatorname{Hom}(W,W_{\hbar}((x)))\subset(\textnormal{{End}}W)[[x,x^{-1}]].$
Note that ${\mathcal{E}}_{\hbar}(W)={\mathcal{E}}(W_{0})[[\hbar]]$ is a topologically free $\mathbb{C}$$\hbar$]]-module. On the other hand, in terms of the natural $\mathbb{C}$$\hbar$]]-module maps
(3.4) $\pi_{n}:\ (\textnormal{{End}}W)[[x,x^{-1}]]\longrightarrow(\textnormal{{End}}(W/\hbar^{n}W))[[x,x^{-1}]]$
for $n\in{\mathbb{Z}}_{+}$, we have
(3.5) $\quad\quad\quad{\mathcal{E}}_{\hbar}(W)=\{\psi(x)\in(\textnormal{{End}}W)[[x,x^{-1}]]\ |\ \pi_{n}(\psi(x))\in{\mathcal{E}}(W/\hbar^{n}W)\ \ \text{ for }n\geq 1\}.$
It is clear that we have a natural inverse system
$0\longleftarrow W/\hbar W\longleftarrow W/\hbar^{2}W\longleftarrow\cdots\longleftarrow,$
which induces an inverse system
$0\longleftarrow{\mathcal{E}}(W/\hbar W)\longleftarrow{\mathcal{E}}(W/\hbar^{2}W)\longleftarrow\cdots\longleftarrow.$
Then
[…]
3.2. $\hbar$-adic nonlocal vertex algebras and modules
The following notion was introduced in [Li5]:
Definition 3.9.
An $\hbar$-adic nonlocal vertex algebra is a topologically free $\mathbb{C}$$\hbar$]]-module $V$, equipped with a $\mathbb{C}$$\hbar$]]-module map $Y(\cdot,x):V\to(\operatorname{End}V)[[x,x^{-1}]]$ and a distinguished vector ${{\bf 1}}\in V$, satisfying the conditions that
$Y(u,x)v\in V_{\hbar}((x))\quad\text{ for }u,v\in V,$
$Y({\bf 1},x)=1_{V},\quad Y(v,x){\bf 1}\in V[[x]]\ \ \text{ and }\ \ (Y(v,x){\bf 1})|_{x=0}=v\ \text{ for }v\in V,$
and that for any $u,v,w\in V$ and for any $n\in{\mathbb{Z}}_{+}$, there exists $l\in{\mathbb{N}}$ such that
(3.7) $(x_{0}+x_{2})^{l}Y(u,x_{0}+x_{2})Y(v,x_{2})w\equiv(x_{2}+x_{0})^{l}Y(Y(u,x_{0})v,x_{2})w$
modulo $\hbar^{n}V[[x_{0}^{\pm 1},x_{2}^{\pm 1}]]$.
We have (see [Li5]):
Proposition 3.10.
Let $V$ be a topologically free $\mathbb{C}$$\hbar$]]-module equipped with a $\mathbb{C}$$\hbar$]]-module map $Y(\cdot,x):V\to(\operatorname{End}V)[[x,x^{-1}]]$ and a vector ${{\bf 1}}\in V$. Then $(V,Y,{\bf 1})$ is an $\hbar$-adic nonlocal vertex algebra if and only if for every positive integer $n$, $(V/\hbar^{n}V,\overline{Y}_{n},{{\bf 1}})$ is a nonlocal vertex algebra over $\mathbb{C}$}, where $\overline{Y}_{n}(\cdot,x):V/\hbar^{n}V\to(\operatorname{End}(V/\hbar^{n}V))[[x,x^{-1}]]$ is the canonical map induced from $Y(\cdot,x)$.
For any $\hbar$-adic nonlocal vertex algebra $V=V_{0}[[\hbar]]$, define a $\mathbb{C}$$\hbar$]]-module map
$Y(x_{1},x_{2}):\ V\widehat{\otimes}V\widehat{\otimes}{\mathbb{C}}((x))[[\hbar]]\rightarrow(\operatorname{End}V)[[x_{1}^{\pm 1},x_{2}^{\pm}]]$
by
(3.8) $Y(x_{1},x_{2})(u\otimes v\otimes f(x))=f(x_{1}-x_{2})Y(u,x_{1})Y(v,x_{2})$
for $u,v\in V_{0},\ f(x)\in{\mathbb{C}}((x))$.
Definition 3.11.
An $\hbar$-adic weak quantum vertex algebra is an $\hbar$-adic nonlocal vertex algebra $V$ which satisfies the ($\hbar$-adic) $\mathcal{S}$}-locality: For any $u,v\in V$, there exists $F(u,v,x)\in V\widehat{\otimes}V\widehat{\otimes}{\mathbb{C}}((x))[[\hbar]]$, satisfying the condition that for any $n\in{\mathbb{Z}}_{+}$, there is $k\in{\mathbb{N}}$ such that
(3.9) $(x_{1}-x_{2})^{k}Y(u,x_{1})Y(v,x_{2})\equiv(x_{1}-x_{2})^{k}Y(x_{2},x_{1})(F(u,v,x))$
modulo $\hbar^{n}(\operatorname{End}V)[[x_{1}^{\pm 1},x_{2}^{\pm 1}]]$.
Definition 3.12.
Let $M$ be a $\mathbb{C}$$\hbar$]]-module. For $A(x_{1},x_{2}),B(x_{1},x_{2})\in M[[x_{1}^{\pm 1},x_{2}^{\pm 1}]]$, we write $A(x_{1},x_{2})\sim B(x_{1},x_{2})$ if for any $n\in{\mathbb{Z}}_{+}$, there exists $k\in{\mathbb{N}}$ such that
(3.10) $(x_{1}-x_{2})^{k}A(x_{1},x_{2})\equiv(x_{1}-x_{2})^{k}B(x_{1},x_{2})\ \mod\hbar^{n}M[[x_{1}^{\pm 1},x_{2}^{\pm 1}]].$
Let $V$ be an $\hbar$-adic weak quantum vertex algebra. From definition, there exists a $\mathbb{C}$$\hbar$]]-module map $F(x):V\widehat{\otimes}V\rightarrow V\widehat{\otimes}V\widehat{\otimes}{\mathbb{C}}((x))[[\hbar]]$ such that
(3.11) $Y(u,x_{1})Y(v,x_{2})\sim Y(x_{2},x_{1})(F(x)(v\otimes u))\quad\text{ for }u,v\in V.$
As before, such a $\mathbb{C}$$\hbar$]]-module map $F(x)$ is called an $\mathcal{S}$}-locality operator of $V$.
We have (see [Li5, (2.25), Prop. 2.19]):
Proposition 3.13.
Let $V$ be an $\hbar$-adic weak quantum vertex algebra and let ${\mathcal{S}}(x)$ be an $\mathcal{S}$}-locality operator of $V$. Then
(3.12) $x_{0}^{-1}\delta\!\left(\frac{x_{1}-x_{2}}{x_{0}}\right)\!Y(u,x_{1})Y(v,x_{2})w$
$\quad-x_{0}^{-1}\delta\!\left(\frac{x_{2}-x_{1}}{-x_{0}}\right)\!Y(x_{2},x_{1})({\mathcal{S}}(x)(v\otimes u))w$
[…]
for $u,v,w\in V$, which is called the $\mathcal{S}$}-Jacobi identity for the ordered triple $(u,v,w)$.
As an immediate consequence of the $\mathcal{S}$}-Jacobi identity, we have
(3.13) $Y(u,x_{1})Y(v,x_{2})-Y(x_{2},x_{1})(\mathcal{S}(x)(v\otimes u))$
[…]
(the $\mathcal{S}$-commutator formula) for $u,v\in V$.
For an $\hbar$-adic nonlocal vertex algebra $V$, define a $\mathbb{C}$$\hbar$]]-linear operator $\mathcal{D}$ on $V$ just as for a nonlocal vertex algebra (see (2.9)). Then the properties (2.10)–(2.12) hold.
Let $W$ be a topologically free $\mathbb{C}$$\hbar$]]-module. A rational quantum Yang-Baxter operator on $W$ is a $\mathbb{C}$$\hbar$]]-module map
(3.14) ${\mathcal{S}}(x):\ W\widehat{\otimes}W\to W\widehat{\otimes}W\widehat{\otimes}{\mathbb{C}}((x))[[\hbar]],$
satisfying the rational quantum Yang-Baxter equation.
The following notion, formulated in [Li5], slightly generalizes Etingof-Kazhdan’s notion of quantum vertex operator algebra (see [EK]; cf. Remark 3.15 below):
Definition 3.14.
An $\hbar$-adic quantum vertex algebra is an $\hbar$-adic nonlocal vertex algebra $V$ equipped with a unitary rational quantum Yang-Baxter operator ${\mathcal{S}}(x)$ on $V$, such that ${\mathcal{S}}(x)$ is an $\mathcal{S}$}-locality operator and such that the shift condition (2.20) and the hexagon identity (2.21) on $V\widehat{\otimes}V\widehat{\otimes}V$ hold.
Remark 3.15.
Note that a quantum vertex operator algebra in the sense of [EK] is an $\hbar$-adic quantum vertex algebra $V$ such that $V/\hbar V$ is a vertex algebra. For a general $\hbar$-adic quantum vertex algebra $V$, $V/\hbar V$ is a quantum vertex algebra.
The following is a reformulation of Proposition 1.11 of [EK]:
Proposition 3.16.
Let $V$ be an $\hbar$-adic weak quantum vertex algebra which is non-degenerate in the sense that the nonlocal vertex algebra $V/\hbar V$ is non-degenerate. Then there exists a unique $\mathcal{S}$}-locality operator ${\mathcal{S}}(x)$. Furthermore, ${\mathcal{S}}(x)$ is a unitary rational quantum Yang-Baxter operator and $(V,{\mathcal{S}}(x))$ is an $\hbar$-adic quantum vertex algebra.
Let $V$ be an $\hbar$-adic nonlocal vertex algebra and let $U$ be a subset. Set $U^{(1)}={\mathbb{C}}[[\hbar]](U\cup\{{\bf 1}\})$ and then inductively define $U^{(n+1)}$ for $n\geq 1$ by
(3.15) $U^{(n+1)}={\rm span}_{{\mathbb{C}}[[\hbar]]}\{a_{m}b\ |\ a,b\in U^{(n)},\ m\in{\mathbb{Z}}\}.$
We have ${\bf 1}\in U^{(n)}$ and $U^{(n)}\subset U^{(n+1)}$ for all $n\in{\mathbb{Z}}_{+}$. Set $U^{(\infty)}=\cup_{n\geq 1}U^{(n)}$. Then $U^{(\infty)}$ is $Y$-closed and contains $U\cup\{{\bf 1}\}$. Furthermore, $\overline{[U^{(\infty)}]}$ is an $\hbar$-adic nonlocal vertex subalgebra, which is called the $\hbar$-adic nonlocal vertex subalgebra generated by $U$.
The following are some straightforward facts:
Lemma 3.17.
Let $U$ be a subset of an $\hbar$-adic nonlocal vertex algebra $V$. Set
(3.16) $U^{\infty}={\rm span}_{{\mathbb{C}}[[\hbar]]}\{u^{(1)}_{m_{1}}\cdots u^{(r)}_{m_{r}}{\bf 1}\ |\ r\geq 0,\ u^{(i)}\in U,\ m_{i}\in{\mathbb{Z}}\}.$
Then
(3.17) $U^{\infty}\subset U^{(\infty)},\quad U^{(\infty)}\subset\overline{U^{\infty}},\quad\overline{U^{(\infty)}}=\overline{U^{\infty}},\quad\overline{[U^{(\infty)}]}=\overline{[U^{\infty}]}.$
In particular, if $V=\overline{[U^{\infty}]}$, then $V=\overline{[U^{(\infty)}]}={\langle}U{\rangle}$.
A subset $U$ of an $\hbar$-adic nonlocal vertex algebra $V$ is called an $\mathcal{S}$}-local subset if for any $u,v\in U$, there exists $A(x)\in({\mathbb{C}}U\otimes{\mathbb{C}}U\otimes{\mathbb{C}}((x)))[[\hbar]]$ such that
$Y(u,x_{1})Y(v,x_{2})\sim Y(x_{2},x_{1})(A(x)).$
We have (see [Li5, Prop. 3.12]):
Lemma 3.18.
Let $V$ be an $\hbar$-adic nonlocal vertex algebra. Suppose $V=U^{(\infty)}[[\hbar]]^{\prime}$ for some $\mathcal{S}$}-local subset $U$. Then $V$ is an $\hbar$-adic weak quantum vertex algebra.
Definition 3.19.
Let be an $\hbar$-adic nonlocal vertex algebra. A $V$-module is a topologically free $\mathbb{C}$$\hbar$]]-module $W$ equipped with a $\mathbb{C}$$\hbar$]]-module map
$Y_{W}(\cdot,x):\ V\rightarrow(\operatorname{End}W)[[x,x^{-1}]],$
satisfying the following conditions:
$Y_{W}(v,x)\in{\mathcal{E}}_{\hbar}(W)\quad\text{ for }v\in V,$
$Y_{W}({\bf 1},x)=1_{W}$, and the ($\hbar$-adic) weak associativity holds: For any $u,v\in V,\ w\in W$ and for every positive integer $n$, there exists $l\in{\mathbb{N}}$ such that
(3.18) $(x_{0}+x_{2})^{l}Y_{W}(u,x_{0}+x_{2})Y_{W}(v,x_{2})w\equiv(x_{2}+x_{0})^{l}Y_{W}(Y(u,x_{0})v,x_{2})w$
modulo $\hbar^{n}W[[x_{0}^{\pm 1},x_{2}^{\pm 1}]]$.
From definition we immediately have:
Proposition 3.20.
Let $V$ be an $\hbar$-adic nonlocal vertex algebra and let $W$ be a topologically free $\mathbb{C}$$\hbar$]]-module equipped with a $\mathbb{C}$$\hbar$]]-module map
$Y_{W}(\cdot,x):\ V\rightarrow(\operatorname{End}W)[[x,x^{-1}]].$
Then $W$ is a $V$-module if and only if $W/\hbar^{n}W$ is a $(V/\hbar^{n}V)$-module for every $n\in{\mathbb{Z}}_{+}$.
For the rest of this subsection, assume that $V$ is an $\hbar$-adic nonlocal vertex algebra and $(W,Y_{W})$ is a $V$-module. Just as for $V$, define a $\mathbb{C}$$\hbar$]]-module map
(3.19) $Y_{W}(x_{1},x_{2}):\ V\widehat{\otimes}V\widehat{\otimes}{\mathbb{C}}((x))[[\hbar]]\rightarrow(\operatorname{End}W)[[x_{1}^{\pm 1},x_{2}^{\pm 1}]]$
by
$Y_{W}(x_{1},x_{2})(u\otimes v\otimes f(x))=f(x_{1}-x_{2})Y_{W}(u,x_{1})Y_{W}(v,x_{2})$
for $u,v\in V_{0},\ f(x)\in{\mathbb{C}}((x))$, where $V=V_{0}[[\hbar]]$.
The following are some standard results (see [Li5], Prop. 2.25; cf. [Li2]):
Lemma 3.21.
Let
$u,v\in V,\ A(x)\in V\widehat{\otimes}V\widehat{\otimes}{\mathbb{C}}((x))[[\hbar]].$
Then
$Y_{W}(u,x_{1})Y_{W}(v,x_{2})\sim Y_{W}(x_{2},x_{1})(A(x))$
if and only if
$x_{0}^{-1}\delta\!\left(\frac{x_{1}-x_{2}}{x_{0}}\right)\!Y_{W}(u,x_{1})Y_{W}(v,x_{2})-x_{0}^{-1}\delta\!\left(\frac{x_{2}-x_{1}}{-x_{0}}\right)\!Y_{W}(x_{2},x_{1})(A(x))$
$\hskip 56.9055pt=x_{1}^{-1}\delta\!\left(\frac{x_{2}+x_{0}}{x_{1}}\right)\!Y_{W}(Y(u,x_{0})v,x_{2}).$
Proposition 3.22.
Let
$u,v,c^{(0)},c^{(1)},\dots\in V,\ A(x)\in V\widehat{\otimes}V\widehat{\otimes}{\mathbb{C}}((x))[[\hbar]]$
with $\lim_{n\rightarrow\infty}c^{(n)}=0$. If
$Y(u,x_{1})Y(v,x_{2})-Y(x_{2},x_{1})(A(x))=\sum_{n\geq 0}Y(c^{(n)},x_{2})\frac{1}{n!}\!\left(\frac{\partial}{\partial x_{2}}\right)^{n}x_{1}^{-1}\delta\!\left(\frac{x_{2}}{x_{1}}\right)$
on $V$, then
$Y_{W}(u,x_{1})Y_{W}(v,x_{2})-Y_{W}(x_{2},x_{1})(A(x))=\sum_{n\geq 0}Y_{W}(c^{(n)},x_{2})\frac{1}{n!}\!\left(\frac{\partial}{\partial x_{2}}\right)^{n}x_{1}^{-1}\delta\!\left(\frac{x_{2}}{x_{1}}\right)\!.$
Furthermore, if $(W,Y_{W})$ is faithful, the converse is also true with $u_{n}v=c^{(n)}$ for $n\geq 0$.
Proposition 3.23.
Let
$u,v\in V,\ A(x)\in V\widehat{\otimes}V\widehat{\otimes}{\mathbb{C}}((x))[[\hbar]].$
If
$Y(u,x_{1})Y(v,x_{2})\sim Y(x_{2},x_{1})(A(x))$
then
$Y_{W}(u,x_{1})Y_{W}(v,x_{2})\sim Y_{W}(x_{2},x_{1})(A(x)).$
The converse is also true if $(W,Y_{W})$ is a faithful module.
We also have (see [Li5, Prop. 3.13]):
Lemma 3.24.
Let $e\in W$ and let $U$ be an $\mathcal{S}$}-local subset of $V$ such that
$Y_{W}(u,x)e\in W[[x]]\quad\text{ for }u\in U.$
Then $Y_{W}(v,x)e\in W[[x]]$ for all $v\in{\langle}U{\rangle}.$ Furthermore, the map $\theta:\ {\langle}U{\rangle}\rightarrow W,$ defined by $\theta(v)=v_{-1}e$ for $v\in{\langle}U{\rangle}$, is a ${\langle}U{\rangle}$-module homomorphism.
We present the following two nonstandard results for studying double Yangians:
Lemma 3.25.
Let $a,b,u,v\in V,\ R(x),f(x),g(x)\in{\mathbb{C}}((x))[[\hbar]].$ If
(3.20) $Y(a,x)Y(b,z)-R(z-x)Y(b,z)Y(a,x)=f(x-z)Y(u,z)+g(z-x)Y(v,z)$
on $V$, then
(3.21) $\ \quad Y_{W}(a,x)Y_{W}(b,z)-R(z-x)Y_{W}(b,z)Y_{W}(a,x)$
$=f(x-z)Y_{W}(u,z)+g(z-x)Y_{W}(v,z)$
on $W$. Furthermore, the converse is also true if $(W,Y_{W})$ is faithful.
Proof.
Write $f(x)=\sum_{m\in{\mathbb{Z}}}f(m)x^{m}$ with $f(m)\in{\mathbb{C}}[[\hbar]]$ and define
(3.22) $\tilde{f}(z)=\operatorname{Res}_{x}f(x)e^{zx}=\sum_{n\geq 0}\frac{1}{n!}f(-n-1)z^{n}\in{\mathbb{C}}[[\hbar]][[z]].$
Note that
$f(x-z)-f(-z+x)=\tilde{f}\!\left(\frac{\partial}{\partial z}\right)x^{-1}\delta\!\left(\frac{z}{x}\right)\!.$
Using this we rewrite (3.20) as
(3.23) $Y(a,x)Y(b,z)$
$\ \ -\left(R(z-x)Y(b,z)Y(a,x)+f(-z+x)Y(u,z)+g(z-x)Y(v,z)\right)$
$=\$$(f(x-z)-f(-z+x))Y(u,z)$
[…]
Similarly, we rewrite (3.21) as
$Y_{W}(a,x)Y_{W}(b,z)$
$\ \ -\left(R(z-x)Y_{W}(b,z)Y_{W}(a,x)+f(-z+x)Y_{W}(u,z)+g(z-x)Y_{W}(v,z)\right)$
[…]
Note that $\lim_{n\rightarrow\infty}f(-n)=0$ as $f(x)\in{\mathbb{C}}((x))[[\hbar]]$. Then it follows immediately from Propositions 3.22. ∎
Lemma 3.26.
Let
$u,v\in V,\ B(x)\in V\widehat{\otimes}V\widehat{\otimes}{\mathbb{C}}((x))[[\hbar]],\ p(x,z),q(x,z)\in{\mathbb{C}}[x,z]$
with $p(x,0)\neq 0$. If
(3.24) $p(x_{1}-x_{2},\hbar)Y(u,x_{1})Y(v,x_{2})=q(x_{2}-x_{1},\hbar)Y(x_{2},x_{1})(B(x))$
on $V$, then
(3.25) $p(x_{1}-x_{2},\hbar)Y_{W}(u,x_{1})Y_{W}(v,x_{2})=q(x_{2}-x_{1},\hbar)Y_{W}(x_{2},x_{1})(B(x))$
on $W$. The converse is also true if $(W,Y_{W})$ is a faithful $V$-module.
Proof.
Rewrite (3.24) as
$p(x_{1}-x_{2},\hbar)Y(u,x_{1})Y(v,x_{2})=p(x_{1}-x_{2},\hbar)\left(\frac{q(x_{2}-x_{1},\hbar)}{p(-x_{2}+x_{1},\hbar)}\right)Y(x_{2},x_{1})(B(x)),$
where as a convention,
$\frac{q(x_{2}-x_{1},\hbar)}{p(-x_{2}+x_{1},\hbar)}:=e^{-x_{1}\frac{\partial}{\partial x_{2}}}\iota_{x_{2},\hbar}\left(\frac{q(x_{2},\hbar)}{p(-x_{2},\hbar)}\right)\!.$
The same argument in [Li5, Lemma 4.25] proves
(3.26) $Y(u,x_{1})Y(v,x_{2})\sim\frac{q(x_{2}-x_{1},\hbar)}{p(-x_{2}+x_{1},\hbar)}Y(x_{2},x_{1})(B(x)).$
Then by Lemma 3.21 we have
(3.27) $x_{0}^{-1}\delta\!\left(\frac{x_{1}-x_{2}}{x_{0}}\right)\!Y(u,x_{1})Y(v,x_{2})$
$\quad-x_{0}^{-1}\delta\!\left(\frac{x_{2}-x_{1}}{-x_{0}}\right)\frac{q(x_{2}-x_{1},\hbar)}{p(-x_{2}+x_{1},\hbar)}Y(x_{2},x_{1})(B(x))$
[…]
Using this and (3.24) we get
$\operatorname{Res}_{x_{0}}x_{1}^{-1}\delta\!\left(\frac{x_{2}+x_{0}}{x_{1}}\right)p(x_{0},\hbar)Y(Y(u,x_{0})v,x_{2})=0.$
Then
(3.28) $\operatorname{Res}_{x_{0}}x_{1}^{-1}\delta\!\left(\frac{x_{2}+x_{0}}{x_{1}}\right)p(x_{0},\hbar)Y(u,x_{0})v$
[…]
$=\$$0.$
With (3.26), by Proposition 3.23 we have
$Y_{W}(u,x_{1})Y_{W}(v,x_{2})\sim\frac{q(x_{2}-x_{1},\hbar)}{p(-x_{2}+x_{1},\hbar)}Y_{W}(x_{2},x_{1})(B(x)),$
which by Lemma 3.21 is equivalent to
(3.29) $x_{0}^{-1}\delta\!\left(\frac{x_{1}-x_{2}}{x_{0}}\right)\!Y_{W}(u,x_{1})Y_{W}(v,x_{2})$
$\quad-x_{0}^{-1}\delta\!\left(\frac{x_{2}-x_{1}}{-x_{0}}\right)\!\frac{q(x_{2}-x_{1},\hbar)}{p(-x_{2}+x_{1},\hbar)}Y_{W}(x_{2},x_{1})(B(x))$
[…]
Applying $\operatorname{Res}_{x_{0}}p(x_{1}-x_{2},\hbar)$ and then using (3.28) we obtain (3.25) as
$p(x_{1}-x_{2},\hbar)Y_{W}(u,x_{1})Y_{W}(v,x_{2})-Y_{W}(x_{2},x_{1})(B(x))$
[…]
$=\$$0.$
Conversely, assume that $(W,Y_{W})$ is a faithful $V$-module and (3.25) holds. By Proposition 3.23, (3.26) holds, and then by Lemma 3.21, (3.27) holds. On the other hand, with (3.25) we have (3.29). As $(W,Y_{W})$ is faithful, we conclude that (3.28) holds. Then we obtain (3.24). ∎
Definition 3.27.
Let $V$ be an $\hbar$-adic nonlocal vertex algebra. An $\hbar$-adic nonlocal vertex subalgebra of $V$ is a $\mathbb{C}$$\hbar$]]-submodule $U$ such that $U$ itself is $\hbar$-adically complete,
(3.30) ${\bf 1}\in U,\ \ u_{m}v\in U\quad\text{ for }u,v\in U,\ m\in{\mathbb{Z}},$
and $\lim_{m\rightarrow\infty}u_{m}v=0$ with respect to the $\hbar$-adic topology of $U$. Furthermore, $U$ is called a strong $\hbar$-adic nonlocal vertex subalgebra if $U=[U]$ in $V$.
Remark 3.28.
From definition, a $\mathbb{C}$$\hbar$]]-submodule $U$ is an $\hbar$-adic nonlocal vertex subalgebra if and only if there is an $\hbar$-adic nonlocal vertex algebra structure on $U$ such that the natural embedding of $U$ into $V$ is a homomorphism. On the other hand, a $\mathbb{C}$$\hbar$]]-submodule $U$ is a strong $\hbar$-adic nonlocal vertex subalgebra if and only if $[U]=U$, $U$ is $\hbar$-adically complete, and (3.30) holds.
Definition 3.29.
Let $V$ be an $\hbar$-adic nonlocal vertex algebra. A strong ideal of $V$ is a $\mathbb{C}$$\hbar$]]-submodule $J$ such that $[J]=J$, $J$ is $\hbar$-adically complete, and
(3.31) $v_{m}J\subset J,\ \ a_{m}V\subset J\quad\text{ for }v\in V,\ m\in{\mathbb{Z}},\ a\in J.$
An $\hbar$-adic nonlocal vertex algebra $V$ is said to be relatively simple if it is nonzero and if $V$ is its only nonzero strong ideal.
The following is straightforward:
Lemma 3.30.
Let $\psi:V\rightarrow K$ be a homomorphism of $\hbar$-adic nonlocal vertex algebras. Then $\ker(\psi)$ is a strong ideal of $V$.
Using Lemma 3.8 we immediately have:
Lemma 3.31.
Let $J$ be a strong ideal of an $\hbar$-adic nonlocal vertex algebra $V$. Then the $\mathbb{C}$$\hbar$]]-module $V/J$ is topologically free and it is an $\hbar$-adic nonlocal vertex algebra.
Lemma 3.32.
Let $T$ be a subset of an $\hbar$-adic nonlocal vertex algebra. Denote by $(T)$ the intersection of all strong ideals of $V$ containing $T$. Then $(T)$ is the smallest strong ideal containing $T$.
Proposition 3.33.
Let $V$ be an $\hbar$-adic nonlocal vertex algebra such that $V/\hbar V$ is a simple nonlocal vertex algebra over $\mathbb{C}$}. Then $V$ is relatively simple.
Proof.
Let $J$ be any nonzero strong ideal of $V$. In view of Lemma 3.8, there exist a $\mathbb{C}$}-subspace $V_{0}$ of $V$ and a $\mathbb{C}$}-subspace $J_{0}$ of $V_{0}$ such that $V=V_{0}[[\hbar]]$ and $J=J_{0}[[\hbar]]$. Consider the natural $\mathbb{C}$$\hbar$]]-morphism $\theta:J\rightarrow V\rightarrow V/\hbar V.$ We have
$\ker\theta=J\cap\hbar V=\hbar J.$
Then $\theta$ reduces to an injective $\mathbb{C}$$\hbar$]]-morphism $\bar{\theta}:J/\hbar J\rightarrow V/\hbar V$. It is clear that $J/\hbar J$ is an ideal of $V/\hbar V$. Note that since $J\neq 0$, we have $J\neq\hbar J$, i.e., $J/\hbar J\neq 0$. As $V/\hbar V$ is a simple nonlocal vertex algebra (over $\mathbb{C}$}), we conclude that $J/\hbar J=V/\hbar V$, which implies $J_{0}=V_{0}$. Therefore, we have $J=J_{0}[[\hbar]]=V_{0}[[\hbar]]=V$. ∎
Let $W$ be a module for an $\hbar$-adic nonlocal vertex algebra $V$. A strong $V$-submodule of $W$ is a $\mathbb{C}$$\hbar$]]-submodule $U$ such that $[U]=U$, $U$ is $\hbar$-adically complete, and
(3.32) $a_{m}u\subset U\quad\text{ for }a\in V,\ m\in{\mathbb{Z}},\ u\in U.$
A $V$-module $W$ is said to be relatively simple (irreducible) if $W\neq 0$ and if $W$ is its only nonzero strong $V$-submodule.
Using Lemma 3.8, we immediately have:
Lemma 3.34.
Let $W$ be a module for an $\hbar$-adic nonlocal vertex algebra $V$ and let $U$ be a strong $V$-submodule of $W$. Then $W/U$ is topologically free and it is a $V$-module.
On the other hand, from the proof of Lemma 3.33 we immediately have:
Proposition 3.35.
Let $V$ be an $\hbar$-adic nonlocal vertex algebra and let $W$ be a $V$-module such that $W/\hbar W$ is an irreducible $V/\hbar V$-module. Then $W$ is relatively simple.
3.3. General constructions
Let $W=W_{0}[[\hbar]]$ be a topologically free $\mathbb{C}$$\hbar$]]-module. Recall
${\mathcal{E}}_{\hbar}(W)={\mathcal{E}}(W_{0})[[\hbar]]=(\textnormal{{Hom}}(W_{0},W_{0}((x))))[[\hbar]]$
and for each positive integer $n$, we have a natural linear map
(3.33) $\pi_{n}:\ (\operatorname{End}W)[[x,x^{-1}]]\longrightarrow(\operatorname{End}(W/\hbar^{n}W))[[x,x^{-1}]].$
Let $a(x)\in(\operatorname{End}W)[[x,x^{-1}]]$. Then $a(x)\in{\mathcal{E}}_{\hbar}(W)$ if and only if $\pi_{n}(a(x))\in{\mathcal{E}}(W/\hbar^{n}W)$ for every positive integer $n$.
Definition 3.36.
A finite sequence in ${\mathcal{E}}_{\hbar}(W)$ is said to be $\hbar$-adically compatible if it is compatible in ${\mathcal{E}}(W/\hbar^{n}W)$ for every positive integer $n$. A subset $U$ of ${\mathcal{E}}_{\hbar}(W)$ is said to be $\hbar$-adically compatible if every finite sequence in $U$ is $\hbar$-adically compatible, i.e., for every positive integer $n$, $\pi_{n}(U)$ is a compatible subset of ${\mathcal{E}}(W/\hbar^{n}W)$.
Let $(a(x),b(x))$ be an $\hbar$-adically compatible pair in ${\mathcal{E}}_{\hbar}(W)$. For every positive integer $n$, as $(\pi_{n}(a(x)),\pi_{n}(b(x))$ is a compatible pair in ${\mathcal{E}}(W/\hbar^{n}W)$, we have $\pi_{n}(a(x))_{m}\pi_{n}(b(x))\in{\mathcal{E}}(W/\hbar^{n}W)$ for $m\in{\mathbb{Z}}$. For $m\in{\mathbb{Z}}$, define $a(x)_{m}b(x)\in{\mathcal{E}}_{\hbar}(W)$ by
[…]
Form a generating function
(3.35) $Y_{{\mathcal{E}}}(a(x),z)b(x)=\sum_{m\in{\mathbb{Z}}}a(x)_{m}b(x)z^{-m-1}.$
An $\hbar$-adically compatible $\mathbb{C}$$\hbar$]]-submodule $K$ of ${\mathcal{E}}_{\hbar}(W)$ is said to be $Y_{{\mathcal{E}}}$-closed if
[…]
The following result was obtained in [Li5, Theorem 4.16]:
Theorem 3.37.
Let $W$ be a topologically free $\mathbb{C}$$\hbar$]]-module. Assume that $V$ is an $\hbar$-adically compatible $\mathbb{C}$$\hbar$]]-submodule of ${\mathcal{E}}_{\hbar}(W)$ such that $V$ contains $1_{W}$, $V$ is $Y_{\mathcal{E}}$-closed, $[V]=V$, and $V$ is $\hbar$-adically complete. Then $(V,Y_{\mathcal{E}},1_{W})$ carries the structure of an $\hbar$-adic nonlocal vertex algebra and $W$ is a faithful $V$-module with $Y_{W}(a(x),z)=a(z)$ for $a(x)\in V$. On the other hand, let $U$ be any $\hbar$-adically compatible subset of ${\mathcal{E}}_{\hbar}(W)$. Then there exists a (unique) smallest $\hbar$-adically compatible $\mathbb{C}$$\hbar$]]-submodule ${\langle}U{\rangle}$ of ${\mathcal{E}}_{\hbar}(W)$ such that ${\langle}U{\rangle}$ contains $\cup\{1_{W}\}, U{\langle}U{\rangle}$ is $Y_{\mathcal{E}}$-closed, ${\langle}U{\rangle}]={\langle}U{\rangle}$, and ${\langle}U{\rangle}$ is $\hbar$-adically complete.
By a straightforward argument we have:
Lemma 3.38.
Let $(W,Y_{W})$ be a module for an $\hbar$-adic nonlocal vertex algebra $V$. For $u,v\in V$, if $(Y_{W}(u,x),Y_{W}(v,x))$ is $\hbar$-adically compatible, then
(3.36) $Y_{W}(Y(u,z)v,x)=Y_{{\mathcal{E}}}(Y_{W}(u,x),z)Y_{W}(v,x).$
Define a $\mathbb{C}$$\hbar$]]-module map
$Z_{W}(x_{1},x_{2}):\ \ {\mathcal{E}}_{\hbar}(W)\widehat{\otimes}{\mathcal{E}}_{\hbar}(W)\widehat{\otimes}{\mathbb{C}}((z))[[\hbar]]\to(\operatorname{End}W)[[x_{1}^{\pm 1},x_{2}^{\pm 1}]]$
by
(3.37) $Z_{W}(x_{1},x_{2})(a(x)\otimes b(x)\otimes f(z))=f(x_{1}-x_{2})a(x_{1})b(x_{2}).$
Definition 3.39.
A subset $U$ of ${\mathcal{E}}_{\hbar}(W)$ is said to be $\mathcal{S}$}-local if for any $a(x),b(x)\in U$, there exists $A(z)\in({\mathbb{C}}U\otimes{\mathbb{C}}U\otimes{\mathbb{C}}((z)))[[\hbar]]$ such that
(3.38) $a(x_{1})b(x_{2})\sim Z_{W}(x_{2},x_{1})(A(z)).$
We have (see [Li5, Lemmas 4.23]):
Lemma 3.40.
Let $U$ be an $\mathcal{S}$-local subset of ${\mathcal{E}}_{\hbar}(W)$. Then $U$ is $\hbar$-adically compatible.
The following is a slight modification of [Li5, Theorem 4.24] (cf. [Li3, Thm 2.9]):
Theorem 3.41.
Let $V$ be a topologically free $\mathbb{C}$$\hbar$]]-module, $U$ a subset, $\bf 1$} a vector of $V$, and $Y_{0}(\cdot,x)$ a map from $U$ to ${\mathcal{E}}_{\hbar}(V)$. Assume all the following conditions hold:
(3.39) $Y_{0}(u,x){\bf 1}\in V[[x]]\ \text{ and }\ \lim_{x\rightarrow 0}Y_{0}(u,x){\bf 1}=u\ \text{ for }u\in U,$
$U(x):=\!\{u(x):=Y_{0}(u,x)=\sum_{n\in{\mathbb{Z}}}u_{n}x^{-n-1}|u\in U\}$ is $\mathcal{S}$}-local, and $V=\overline{[U^{\infty}]}$, where
(3.40) $U^{\infty}={\rm span}_{{\mathbb{C}}[[\hbar]]}\{u^{(1)}_{n_{1}}\cdots u^{(r)}_{n_{r}}{\bf 1}\ |\ r\geq 0,\ u^{(i)}\in U,\ n_{i}\in{\mathbb{Z}}\}.$
In addition, assume that there exists a $\mathbb{C}$$\hbar$]]-module map $\psi_{x}:V\rightarrow{\langle}U(x){\rangle}$ such that $\psi_{x}({\bf 1})=1_{V}$ and
(3.41) $\psi_{x}(u_{n}v)=u(x)_{n}\psi_{x}(v)\quad\text{ for }u\in U,\ n\in{\mathbb{Z}},\ v\in V.$
Then $Y_{0}(\cdot,x)$ can be extended uniquely to a $\mathbb{C}$$\hbar$]]-module map $Y(\cdot,x):V\rightarrow{\mathcal{E}}_{\hbar}(V)$ such that $(V,Y,{\bf 1})$ carries the structure of an $\hbar$-adic nonlocal vertex algebra and $\psi_{x}$ is an isomorphism of $\hbar$-adic nonlocal vertex algebras. Furthermore, if $V=U^{\infty}[[\hbar]]^{\prime}$, then $V$ is an $\hbar$-adic weak quantum vertex algebra.
Proof.
First, with $U(x)$ an $\mathcal{S}$}-local subset of ${\mathcal{E}}_{\hbar}(V)$ we have an $\hbar$-adic nonlocal vertex algebra ${\langle}U(x){\rangle}$ with $V$ as a module. Second, by Lemma 3.24 we get $a(x){\bf 1}\in V[[x]]$ for $a(x)\in{\langle}U(x){\rangle}$ and the map $\phi:{\langle}U(x){\rangle}\rightarrow V$, defined by $\phi(a(x))=(a(x){\bf 1})|_{x=0}$ for $a(x)\in{\langle}U(x){\rangle}$ is a ${\langle}U(x){\rangle}$-module morphism with $\phi(1_{V})={\bf 1}$. Notice that
$\psi_{x}(u)=\psi_{x}(u_{-1}{\bf 1})=u(x)_{-1}1_{V}=u(x)\ \text{ and }\ \phi(u(x))=\lim_{x\rightarrow 0}u(x){\bf 1}=u$
for $u\in U$. It follows that
$\phi\circ\psi_{x}=1\ \text{ on }U^{\infty}\ \text{ and }\ \psi_{x}\circ\phi=1\ \text{ on }U(x)^{\infty}.$
Then $\phi\circ\psi_{x}=1\ \text{ on }\overline{[U^{\infty}]}=V.$ On the other hand, we have $\psi_{x}\circ\phi=1$ on $\overline{U(x)^{\infty}}$. As $U(x)^{(\infty)}\subset\overline{U(x)^{\infty}}$ by Lemma 3.17, we have $\psi_{x}\circ\phi=1$ on $U(x)^{(\infty)}$, which implies $\psi_{x}\circ\phi=1$ on $\overline{[U(x)^{(\infty)}]}={\langle}U(x){\rangle}$. Thus, $\psi_{x}$ and $\phi$ are inverses each other. The rest follows from the proof of [Li5, Theorem 4.24]. ∎
The following is a straightforward $\hbar$-adic analogue of ([Li3, Theorem 2.11]):
Theorem 3.42.
Let $V$ be an $\hbar$-adic weak quantum vertex algebra, let $U$ be a subset of $V$ such that $V={\langle}U{\rangle}$, let $W$ be a topologically free $\mathbb{C}$$\hbar$]]-module, and let $Y_{W}^{0}(\cdot,x):U\rightarrow{\mathcal{E}}_{\hbar}(W)$ be a map. Suppose that $Y_{W}^{0}(U)$ is $\mathcal{S}$}-local and that there exists a $\mathbb{C}$$\hbar$]]-module map $\psi_{x}:V\rightarrow{\langle}Y_{W}^{0}(U){\rangle}$ such that $\psi_{x}({\bf 1})=1_{W}$ and
(3.42) $\psi_{x}(u_{n}v)=u(x)_{n}\psi_{x}(v)\quad\text{ for }u\in U,\ n\in{\mathbb{Z}},\ v\in V,$
where $u(x):=Y_{W}^{0}(u,x)$. Then $Y_{W}^{0}(\cdot,x)$ can be extended uniquely to a $\mathbb{C}$$\hbar$]]-module map $Y_{W}(\cdot,x):V\rightarrow{\mathcal{E}}_{\hbar}(W)$ such that $(W,Y_{W})$ carries the structure of a $V$-module.
To conclude this subsection, we present some technical results which we need. Let $W$ be a topologically free $\mathbb{C}$$\hbar$]]-module. The following is a straightforward $\hbar$-adic version of a standard result:
Lemma 3.43.
Let $a(x),b(x)\in{\mathcal{E}}_{\hbar}(W)$, $A(z)\in{\mathcal{E}}_{\hbar}(W)\widehat{\otimes}{\mathcal{E}}_{\hbar}(W)\widehat{\otimes}{\mathbb{C}}((z)))[[\hbar]]$ such that
(3.43) $a(x_{1})b(x_{2})\sim Z_{W}(x_{2},x_{1})(A(z)).$
Then $(a(x),b(x))$ is $\hbar$-adically compatible and
(3.44) $x_{0}^{-1}\delta\!\left(\frac{x_{1}-x}{x_{0}}\right)a(x_{1})b(x)-x_{0}^{-1}\delta\!\left(\frac{x-x_{1}}{-x_{0}}\right)\!Z_{W}(x,x_{1})(A(z))$
$\quad\quad=x_{1}^{-1}\delta\!\left(\frac{x+x_{0}}{x_{1}}\right)\!Y_{\mathcal{E}}(a(x),x_{0})b(x).$
Lemma 3.44.
Assume
$a(x),b(x)\in{\mathcal{E}}_{\hbar}(W),\ A(z)\in{\mathcal{E}}_{\hbar}(W)\widehat{\otimes}{\mathcal{E}}_{\hbar}(W)\widehat{\otimes}{\mathbb{C}}((z))[[\hbar]],\ \ p(x,y)\in{\mathbb{C}}(x,y)$
with $p(x,y)\neq 0$ and $\iota_{x,y}(p(x,y)^{\pm 1})\in{\mathbb{C}}((x))[[y]]$ (analytic at $y=0$) such that
(3.45) $p(x_{1}-x_{2},\hbar)a(x_{1})b(x_{2})=Z_{W}(x_{2},x_{1})(A(z)),$
where $p(x_{1}-x_{2},\hbar):=(\iota_{x,y}p(x,y))|_{x=x_{1}-x_{2},y=\hbar}$. Then
(3.46) $a(x_{1})b(x_{2})\sim p(-x_{2}+x_{1},\hbar)^{-1}Z_{W}(x_{2},x_{1})(A(z)).$
Proof.
Let $n$ be any positive integer. Write $\iota_{x,y}(p(x,y)^{-1})=\sum_{i\geq 0}p_{i}^{-}(x)y^{i}$ with $p_{i}^{-}(x)\in{\mathbb{C}}((x))$. Let $k\in{\mathbb{N}}$ such that $x^{k}p_{i}^{-}(x)\in{\mathbb{C}}[[x]]$ for all $0\leq i\leq n-1$. Then we obtain
$(x_{1}-x_{2})^{k}a(x_{1})b(x_{2})$
[…]
$\equiv\$$\sum_{0\leq i<n}(x_{1}-x_{2})^{k}p_{i}^{-}(x_{1}-x_{2})\hbar^{i}\left(p(x_{1}-x_{2},\hbar)a(x_{1})b(x_{2})\right)\ \mod\hbar^{n}$
$=\$$\sum_{0\leq i<n}\left((-x_{2}+x_{1})^{k}p_{i}^{-}(-x_{2}+x_{1})\right)\hbar^{i}\left(p(x_{1}-x_{2},\hbar)a(x_{1})b(x_{2})\right)$
$=\$$\sum_{0\leq i<n}(-x_{2}+x_{1})^{k}p_{i}^{-}(-x_{2}+x_{1})\hbar^{i}\left(Z_{W}(x_{2},x_{1})A(z)\right)$
[…]
$=\$$(x_{1}-x_{2})^{k}p(-x_{2}+x_{1},\hbar)^{-1}Z_{W}(x_{2},x_{1})(A(z)),$
as desired. ∎
Using Lemma 3.44 we immediately have:
Lemma 3.45.
Let $U$ be a subset of ${\mathcal{E}}_{\hbar}(W)$. Assume that for $a(x),b(x)\in U$, there exist
$A(z)\in({\mathbb{C}}U\otimes{\mathbb{C}}U\otimes{\mathbb{C}}((z)))[[\hbar]],\ \ p(x,y)\in{\mathbb{C}}(x,y)$
with $p(x,y)\neq 0$ and $\iota_{x,y}(p(x,y)^{\pm 1})\in{\mathbb{C}}((x))[[y]]$ such that
(3.47) $p(x_{1}-x_{2},\hbar)a(x_{1})b(x_{2})=Z_{W}(x_{2},x_{1})(A(z)).$
Then $U$ is $\mathcal{S}$}-local.
Using Lemmas 3.25 and 3.26, we immediately have:
Lemma 3.46.
Let $U$ be an $\hbar$-adically compatible subset of ${\mathcal{E}}_{\hbar}(W)$, and let
$a(z),b(z)\in U,\ R(x),f(x),g(x)\in{\mathbb{C}}((x))[[\hbar]],\ C(z),D(z)\in{\mathcal{E}}_{\hbar}(W)$
such that
(3.48) $a(z)b(w)-R(w-z)b(w)a(z)=f(z-w)C(w)+g(w-z)D(w).$
Then $C(z),D(z)\in{\langle}U{\rangle}$ and
$Y_{{\mathcal{E}}}(a(x),z)Y_{{\mathcal{E}}}(b(x),w)-R(w-z)Y_{{\mathcal{E}}}(b(x),w)Y_{{\mathcal{E}}}(a(x),z)$
(3.49) $\quad=f(z-w)Y_{{\mathcal{E}}}(C(x),w)+g(w-z)Y_{{\mathcal{E}}}(D(x),w).$
Lemma 3.47.
Let $U$ be an $\hbar$-adically compatible subset of ${\mathcal{E}}_{\hbar}(W)$ with $a(z),b(z)\in U$. Assume
(3.50) $p(z-w,\hbar)a(z)b(w)=q(w-z,\hbar)b(w)a(z),$
where $p(x,z),q(x,z)\in{\mathbb{C}}[x,z]$ with $p(x,0)\neq 0$. Then
(3.51) $p(z-w,\hbar)Y_{{\mathcal{E}}}(a(x),z)Y_{{\mathcal{E}}}(b(x),w)=q(w-z,\hbar)Y_{{\mathcal{E}}}(b(x),w)Y_{{\mathcal{E}}}(a(x),z).$
4. Vertex-operator Serre relation and iterate formula
In this section, we study a Serre-like relation, give two characterizations, and establish a vertex-operator iterate formula.
Lemma 4.1.
Let $U$ be a $\mathbb{C}$$\hbar$]]-module and let $A(z)=\sum_{n\in{\mathbb{Z}}}u(n)z^{-n-1}\in U[[z,z^{-1}]]$, $\mu\in{\mathbb{C}}$. Then $(z-\mu\hbar)A(z)\in U[[z]]$ if and only if ${\rm Sing}_{z}A(z)=u(0)(z-\mu\hbar)^{-1}$.
Proof.
Note that $(z-\mu\hbar)A(z)\in U[[z]]$ if and only if $(z-\mu\hbar){\rm Sing}_{z}A(z)\in U[[z]]$. As
$(z-\mu\hbar){\rm Sing}_{z}A(z)-u(0)=z{\rm Sing}_{z}A(z)-u(0)-\mu\hbar{\rm Sing}_{z}A(z)\in z^{-1}U[[z^{-1}]],$
we see that $(z-\mu\hbar){\rm Sing}_{z}A(z)\in U[[z]]$ implies $(z-\mu\hbar){\rm Sing}_{z}A(z)=u(0)$, which is equivalent to ${\rm Sing}_{z}A(z)=u(0)(z-\mu\hbar)^{-1}$. Conversely, if ${\rm Sing}_{z}A(z)=u(0)(z-\mu\hbar)^{-1}$, then $(z-\mu\hbar){\rm Sing}_{z}A(z)=u(0)$, which implies $(z-\mu\hbar)A(z)\in U[[z]]$. ∎
The following is a straightforward observation (see [JKLT2]):
Lemma 4.2.
Let $b\in{\mathbb{C}},\ F(x,\hbar)\in{\mathbb{C}}[[x,\hbar]]$. Then
(4.1) ${\rm Sing}_{x}(x-b\hbar)^{-1}F(x,\hbar)=(x-b\hbar)^{-1}F(b\hbar,\hbar),$
(4.2) $\operatorname{Res}_{x}(x-b\hbar)^{-1}F(x,\hbar)=F(b\hbar,\hbar).$
From now on, we assume that $W$ is a topologically free $\mathbb{C}$$\hbar$]]-module.
Lemma 4.3.
Let $a(x),b(x)\in{\mathcal{E}}_{\hbar}(W)$. Assume
(4.3) $(x_{1}-x_{2}+\mu\hbar)a(x_{1})b(x_{2})=(x_{1}-x_{2}+\nu\hbar)b(x_{2})a(x_{1}),$
where $\mu,\nu\in{\mathbb{C}}$. Then $(a(x),b(x))$ is $\hbar$-adically compatible and
(4.4) ${\rm Sing}_{z}Y_{{\mathcal{E}}}(a(x),z)b(x)=a(x)_{0}b(x)(z+\mu\hbar)^{-1},$
(4.5) $x_{1}^{-1}\delta\!\left(\frac{x-\mu\hbar}{x_{1}}\right)\!a(x)_{0}b(x)=a(x_{1})b(x)-\!\left(\frac{x-x_{1}-\nu\hbar}{x-x_{1}-\mu\hbar}\right)\!b(x)a(x_{1}).$
Proof.
With the assumption (4.3), by Lemma 3.47 we have
$a(x_{1})b(x_{2})\sim\left(\frac{x_{2}-x_{1}-\nu\hbar}{x_{2}-x_{1}-\mu\hbar}\right)b(x_{2})a(x_{1}),$
which implies that $(a(x),b(x))$ is $\hbar$-adically compatible, and
(4.6) $\operatorname{Res}_{z}x_{1}^{-1}\delta\!\left(\frac{x+z}{x_{1}}\right)\!Y_{{\mathcal{E}}}(a(x),z)b(x)=a(x_{1})b(x)-\left(\frac{x-x_{1}-\nu\hbar}{x-x_{1}-\mu\hbar}\right)b(x)a(x_{1})$
by Lemma 3.43. Then for any nonnegative integer $n$,
$\quad\operatorname{Res}_{z}z^{n}(z+\mu\hbar)Y_{{\mathcal{E}}}(a(x),z)b(x)$
$=\operatorname{Res}_{z}\operatorname{Res}_{x_{1}}(x_{1}-x)^{n}(x_{1}-x+\mu\hbar)x_{1}^{-1}\delta\!\left(\frac{x+z}{x_{1}}\right)\!Y_{{\mathcal{E}}}(a(x),z)b(x)$
$=\operatorname{Res}_{x_{1}}(x_{1}-x)^{n}\left((x_{1}-x+\mu\hbar)a(x_{1})b(x)-(x_{1}-x+\nu\hbar)b(x)a(x_{1})\right)$
$=0.$
This implies that $(z+\mu\hbar)Y_{\mathcal{E}}(a(x),z)b(x)$ is regular in $z$. Then (4.4) holds by Lemma 4.1. Furthermore, using Lemma 4.2 we have
$\operatorname{Res}_{z}x_{1}^{-1}\delta\!\left(\frac{x+z}{x_{1}}\right)\!Y_{{\mathcal{E}}}(a(x),z)b(x)=\$$\operatorname{Res}_{z}x_{1}^{-1}\delta\!\left(\frac{x+z}{x_{1}}\right)\!{\rm Sing}_{z}Y_{{\mathcal{E}}}(a(x),z)b(x)$
[…]
Now, combining this with (4.6) we obtain (4.5). ∎
The following is the Serre relation that we need:
Definition 4.4.
Let $a(x),b(x)\in{\mathcal{E}}_{\hbar}(W)$. The relation
(4.7) $\sum_{\sigma\in S_{2}}\left(a(z_{\sigma(1)})a(z_{\sigma(2)})b(w)-2a(z_{\sigma(1)})b(w)a(z_{\sigma(2)})+b(w)a(z_{\sigma(1)})a(z_{\sigma(2)})\right)=0$
is referred to as the order-$2$ Serre relation for $(a(z_{1}),a(z_{2}),b(w))$.
The following is a key result (cf. [DK, Prop. 1]):
Proposition 4.5.
Let $a(x),b(x)\in{\mathcal{E}}_{\hbar}(W)$ such that
(4.8) $(z-w-2\hbar)a(z)a(w)$ $=(z-w+2\hbar)a(w)a(z),$
(4.9) $(z-w+\hbar)a(z)b(w)$ $=(z-w-\hbar)b(w)a(z).$
Then the order-$2$ Serre relation for $(a(z_{1}),a(z_{2}),b(w))$ holds if and only if
(4.10) $a(z)a(w)_{0}b(w)=a(w)_{0}b(w)a(z)\!\left(\frac{w-z-3\hbar}{w-z-\hbar}\right)\!,$
where
$a(w)_{0}b(w)=\operatorname{Res}_{x_{1}}\!\left(\!a(x_{1})b(w)-\left(\frac{w-x_{1}+\hbar}{w-x_{1}-\hbar}\right)b(w)a(x_{1})\!\right)\!.$
Proof.
We express all the terms in terms of $b(w)a(z_{1})a(z_{2})$. With commutation relation (4.9), by Lemma 4.3 we have
$a(x_{1})b(x)=\left(\frac{x-x_{1}+\hbar}{x-x_{1}-\hbar}\right)b(x)a(x_{1})+x_{1}^{-1}\delta\!\left(\frac{x-\hbar}{x_{1}}\right)\!a(x)_{0}b(x).$
Using this we get
$a(z_{1})a(z_{2})b(w)=\$$a(z_{1})b(w)a(z_{2})\!\left(\frac{w-z_{2}+\hbar}{w-z_{2}-\hbar}\right)\!+a(z_{1})a(w)_{0}b(w)z_{2}^{-1}\delta\!\left(\frac{w-\hbar}{z_{2}}\right)$
[…]
$+a(w)_{0}b(w)a(z_{2})z_{1}^{-1}\delta\!\left(\frac{w-\hbar}{z_{1}}\right)\!\left(\frac{w-z_{2}+\hbar}{w-z_{2}-\hbar}\right)$
$+a(z_{1})a(w)_{0}b(w)z_{2}^{-1}\delta\!\left(\frac{w-\hbar}{z_{2}}\right)$
and
$a(z_{1})b(w)a(z_{2})=b(w)a(z_{1})a(z_{2})\!\left(\frac{w-z_{1}+\hbar}{w-z_{1}-\hbar}\right)\!+a(w)_{0}b(w)a(z_{2})z_{1}^{-1}\delta\!\left(\frac{w-\hbar}{z_{1}}\right)\!.$
Note that
$\frac{w-z_{1}+\hbar}{w-z_{1}-\hbar}\cdot\frac{w-z_{2}+\hbar}{w-z_{2}-\hbar}-2\left(\frac{w-z_{1}+\hbar}{w-z_{1}-\hbar}\right)+1=\frac{2\hbar(z_{2}-z_{1}+2\hbar)}{(w-z_{1}-\hbar)(w-z_{2}-\hbar)},$
$\frac{w-z_{2}+\hbar}{w-z_{2}-\hbar}-2=\frac{z_{2}-w+3\hbar}{w-z_{2}-\hbar}.$
Then we get
(4.11) $a(z_{1})a(z_{2})b(w)-2a(z_{1})b(w)a(z_{2})+b(w)a(z_{1})a(z_{2})$
$=\$$\frac{2\hbar(z_{2}-z_{1}+2\hbar)}{(w-z_{1}-\hbar)(w-z_{2}-\hbar)}b(w)a(z_{1})a(z_{2})$
$+\left(\frac{z_{2}-w+3\hbar}{w-z_{2}-\hbar}\right)a(w)_{0}b(w)a(z_{2})z_{1}^{-1}\delta\!\left(\frac{w-\hbar}{z_{1}}\right)$
$+a(z_{1})a(w)_{0}b(w)z_{2}^{-1}\delta\!\left(\frac{w-\hbar}{z_{2}}\right)\!.$
Noticing that
$2\hbar(z_{2}-z_{1}+2\hbar)b(w)a(z_{1})a(z_{2})+2\hbar(z_{1}-z_{2}+2\hbar)b(w)a(z_{2})a(z_{1})=0,$
we see that the order-$2$ Serre relation for $(a(z_{1}),a(z_{2}),b(w))$ holds if and only if
(4.12) $\left(\left(\frac{z_{2}-w+3\hbar}{w-z_{2}-\hbar}\right)a(w)_{0}b(w)a(z_{2})+a(z_{2})a(w)_{0}b(w)\!\right)\!z_{1}^{-1}\delta\!\left(\frac{w-\hbar}{z_{1}}\right)$
$=\$$-\!\left(\left(\frac{z_{1}-w+3\hbar}{w-z_{1}-\hbar}\right)a(w)_{0}b(w)a(z_{1})+a(z_{1})a(w)_{0}b(w)\!\right)\!z_{2}^{-1}\delta\!\left(\frac{w-\hbar}{z_{2}}\right)\!.$
Notice that multiplying (4.12) by $(w-z_{1}-\hbar)$ and then applying $\operatorname{Res}_{z_{2}}$ gives
(4.13) $(z_{1}-w+\hbar)a(z_{1})a(w)_{0}b(w)=(z_{1}-w+3\hbar)a(w)_{0}b(w)a(z_{1}).$
In view of Lemma 4.3, we have
$a(z_{1})a(w)_{0}b(w)-\!\left(\frac{w-z_{1}-3\hbar}{w-z_{1}-\hbar}\right)\!a(w)_{0}b(w)a(z_{1})=a(w)_{0}a(w)_{0}b(w)z_{1}^{-1}\delta\!\left(\frac{w-\hbar}{z_{1}}\right)\!,$
$a(z_{2})a(w)_{0}b(w)-\!\left(\frac{w-z_{2}-3\hbar}{w-z_{2}-\hbar}\right)\!a(w)_{0}b(w)a(z_{2})=a(w)_{0}a(w)_{0}b(w)z_{2}^{-1}\delta\!\left(\frac{w-\hbar}{z_{2}}\right)\!.$
Combining these relations with (4.12) (applying $\operatorname{Res}_{z_{1}}\operatorname{Res}_{z_{2}}$) we get
$a(w)_{0}a(w)_{0}b(w)=-a(w)_{0}a(w)_{0}b(w),$
which implies $a(w)_{0}a(w)_{0}b(w)=0$. Then we obtain
(4.14) $a(z)a(w)_{0}b(w)-\left(\frac{w-z-3\hbar}{w-z-\hbar}\right)a(w)_{0}b(w)a(z)=0.$
It is clear that this relation also implies (4.12). Now, the proof is complete. ∎
We continue to give another characterization of the Serre relation. First, we present the following two technical results:
Lemma 4.6.
Let $a(x)\in{\mathcal{E}}_{\hbar}(W),\ \mu\in{\mathbb{C}}$ such that
(4.15) $(z-w-\mu\hbar)a(z)a(w)=(z-w+\mu\hbar)a(w)a(z).$
Then
(4.16) $(z-w-\mu\hbar)a(z)a(w)\in{\mathcal{E}}_{\hbar}^{(2)}(W):=\operatorname{Hom}(W,W_{\hbar}((z,w))),$
(4.17) $z^{-1}\delta\left(\frac{w}{z}\right)\left((z-w-\mu\hbar)a(z)a(w)\right)=0.$
Furthermore, we have
(4.18) $(z-w)^{-1}(z-w-\mu\hbar)a(z)a(w)=(w-z)^{-1}(w-z-\mu\hbar)a(w)a(z),$
(4.19) $(z-w)^{-1}(z-w-\mu\hbar)a(z)a(w)\in{\mathcal{E}}_{\hbar}^{(2)}(W).$
Proof.
Set
$A(z,w)=(z-w-\mu\hbar)a(z)a(w).$
From (4.15), we have $A(z,w)\in{\mathcal{E}}_{\hbar}^{(2)}(W)$, so that $A(z,z)$ exists in ${\mathcal{E}}_{\hbar}(W)$. Using (4.15) again we get
$A(z,w)=(z-w+\mu\hbar)a(w)a(z)=-(w-z-\mu\hbar)a(w)a(z)=-A(w,z),$
which implies $A(z,z)=0$. Then
$z^{-1}\delta\!\left(\frac{w}{z}\right)\!\left((z-w-\mu\hbar)a(z)a(w)\right)=z^{-1}\delta\!\left(\frac{w}{z}\right)A(z,w)=z^{-1}\delta\!\left(\frac{w}{z}\right)A(z,z)=0,$
which amounts to
(4.20) $(z-w)^{-1}\left((z-w-\mu\hbar)a(z)a(w)\right)=(-w+z)^{-1}((z-w-\mu\hbar)a(z)a(w)).$
This together with (4.15) gives (4.18), which implies (4.19). ∎
Lemma 4.7.
Let $a(x),b(x)\in{\mathcal{E}}_{\hbar}(W),\ \nu\in{\mathbb{C}}$ such that
(4.21) $(z-w-2\nu\hbar)a(z)a(w)=(z-w+2\nu\hbar)a(w)a(z),$
(4.22) $(z-w+\nu\hbar)a(z)b(w)=(z-w-\nu\hbar)b(w)a(z).$
Set
(4.23) ${\mathopen{\overset{\bullet}{{{\bullet}}}}{a(z)b(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}=(z-w+\nu\hbar)a(z)b(w),$
(4.24) ${\mathopen{\overset{\bullet}{{{\bullet}}}}{a(z_{1})a(z_{2})b(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}$
$=\$$(z_{1}-z_{2})^{-1}(z_{1}-z_{2}-2\nu\hbar)(z_{1}-w+\nu\hbar)(z_{2}-w+\nu\hbar)a(z_{1})a(z_{2})b(w).$
Then
(4.25) ${\mathopen{\overset{\bullet}{{{\bullet}}}}{a(z)b(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}|_{z=w-\nu\hbar}=a(w)_{0}b(w),$
(4.26) $\ z_{2}^{-1}\delta\!\left(\frac{w-\nu\hbar}{z_{2}}\right)\!a(z_{1})\,{\mathopen{\overset{\bullet}{{{\bullet}}}}{a(z_{2})b(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}$
[…]
$\ \ +z_{1}^{-1}\delta\!\left(\frac{w+\nu\hbar}{z_{1}}\right)z_{2}^{-1}\delta\!\left(\frac{w-\nu\hbar}{z_{2}}\right){\mathopen{\overset{\bullet}{{{\bullet}}}}{a(z_{1})a(z_{2})b(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}.$
Proof.
Since
${\mathopen{\overset{\bullet}{{{\bullet}}}}{a(z)b(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}=(z-w+\nu\hbar)a(z)b(w)=(z-w-\nu\hbar)b(w)a(z),$
we have ${\mathopen{\overset{\bullet}{{{\bullet}}}}{a(z)b(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}\in{\mathcal{E}}_{\hbar}^{(2)}(W)$, and
$a(z)b(w)=(z-w+\nu\hbar)^{-1}\ {\mathopen{\overset{\bullet}{{{\bullet}}}}{a(z)b(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}},\quad{\mathopen{\overset{\bullet}{{{\bullet}}}}{a(z)b(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}=-(w-z+\nu\hbar)b(w)a(z).$
Then
$z^{-1}\delta\!\left(\frac{w-\nu\hbar}{z}\right){\mathopen{\overset{\bullet}{{{\bullet}}}}{a(z)b(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}$
$=\$$(z-w+\nu\hbar)^{-1}\ {\mathopen{\overset{\bullet}{{{\bullet}}}}{a(z)b(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}-(-w+z+\nu\hbar)^{-1}\ {\mathopen{\overset{\bullet}{{{\bullet}}}}{a(z)b(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}$
[…]
Applying $\operatorname{Res}_{z}$ and using Lemma 3.43, we get (4.25) as
${\mathopen{\overset{\bullet}{{{\bullet}}}}{a(z)b(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}|_{z=w-\nu\hbar}=\operatorname{Res}_{z}Y_{{\mathcal{E}}}(a(w),z)b(w)=a(w)_{0}b(w).$
Next, we prove (4.26). Using definitions and delta-function substitution we have
$z_{2}^{-1}\delta\!\left(\frac{w-\nu\hbar}{z_{2}}\right)\!a(z_{1})\,{\mathopen{\overset{\bullet}{{{\bullet}}}}{a(z_{2})b(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}$
[…]
and
$z_{2}^{-1}\delta\!\left(\frac{w-\nu\hbar}{z_{2}}\right)\left(\frac{z_{2}-z_{1}-2\nu\hbar}{z_{2}-z_{1}}\right){\mathopen{\overset{\bullet}{{{\bullet}}}}{a(z_{2})b(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}\;a(z_{1})$
[…]
$\times\left((z_{1}-w-\nu\hbar)(z_{2}-z_{1}-2\nu\hbar)(z_{2}-z_{1})^{-1}(z_{2}-w+\nu\hbar)a(z_{2})b(w)a(z_{1})\right)$
[…]
$\times\left((z_{1}-w+\nu\hbar)(z_{1}-z_{2}-2\nu\hbar)(z_{1}-z_{2})^{-1}(z_{2}-w+\nu\hbar)a(z_{1})a(z_{2})b(w)\right)$
[…]
Then we obtain
$z_{2}^{-1}\delta\!\left(\frac{w-\nu\hbar}{z_{2}}\right)\!a(z_{1}){\mathopen{\overset{\bullet}{{{\bullet}}}}{a(z_{2})b(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}-z_{2}^{-1}\delta\!\left(\frac{w-\nu\hbar}{z_{2}}\right)\!\left(\!\frac{z_{2}-z_{1}-2\nu\hbar}{z_{2}-z_{1}}\!\right){\mathopen{\overset{\bullet}{{{\bullet}}}}{a(z_{2})b(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}a(z_{1})$
[…]
as desired. ∎
Now, we present the second characterization of the Serre relation.
Theorem 4.8.
Let $W$ be a topologically free $\mathbb{C}$$\hbar$]]-module and let $a(z),b(z)\in{\mathcal{E}}_{\hbar}(W),\ \nu\in{\mathbb{C}}$ such that (4.21) and (4.22) hold. Then the order-$2$ Serre relation for $(a(z_{1}),a(z_{2}),b(w))$ holds if and only if
(4.27) ${\mathopen{\overset{\bullet}{{{\bullet}}}}{a(z_{1})a(z_{2})b(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}|_{z_{1}=w+\nu\hbar,z_{2}=w-\nu\hbar}=0.$
Proof.
By Proposition 4.5 (with $\hbar$ replaced by $\nu\hbar$), the very Serre relation is equivalent to
(4.28) $a(z_{1})\left(a(w)_{0}b(w)\right)=\left(a(w)_{0}b(w)\right)a(z_{1})\left(\frac{w-z_{1}-3\nu\hbar}{w-z_{1}-\nu\hbar}\right)\!,$
where (recall (4.25))
$a(w)_{0}b(w)=\operatorname{Res}_{x_{1}}\!\left(\!a(x_{1})b(w)-\left(\frac{w-x_{1}+\nu\hbar}{w-x_{1}-\nu\hbar}\right)b(w)a(x_{1})\!\right)\!={\mathopen{\overset{\bullet}{{{\bullet}}}}{a(z)b(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}|_{z=w-\nu\hbar}.$
Applying $\operatorname{Res}_{z_{2}}$ to (4.26) and using delta-function substitution, we get
(4.29) $a(z_{1})\left(a(w)_{0}b(w)\right)-\left(\frac{w-z_{1}-3\nu\hbar}{w-z_{1}-\nu\hbar}\right)\left(a(w)_{0}b(w)\right)a(z_{1})$
[…]
Then we see that (4.28) holds if and only if ${\mathopen{\overset{\bullet}{{{\bullet}}}}{a(z_{1})a(z_{2})b(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}|_{z_{1}=w+\nu\hbar,z_{2}=w-\nu\hbar}=0$. ∎
Next, we consider the situation for $\hbar$-adic nonlocal vertex algebras and modules.
Proposition 4.9.
Let $V$ be an $\hbar$-adic nonlocal vertex algebra with $a,b\in V$ such that
(4.30) $(z-w-2\hbar)Y(a,z)Y(a,w)=(z-w+2\hbar)Y(a,w)Y(a,z),$
(4.31) $(z-w+\hbar)Y(a,z)Y(b,w)=(z-w-\hbar)Y(b,w)Y(a,z).$
Then the order-$2$ Serre relation for $(Y(a,z_{1}),Y(a,z_{2}),Y(b,w))$ holds if and only if
(4.32) ${\rm Sing}_{z}Y(a,z)a_{0}b=0.$
Proof.
From Proposition 4.5 and Lemma 3.43, we see that the order-$2$ Serre relation for $(Y(a,z_{1}),Y(a,z_{2}),Y(b,w))$ holds if and only if
(4.33) $Y(a,z)Y(a_{0}b,w)-\left(\frac{w-z-3\hbar}{w-z-\hbar}\right)\!Y(a_{0}b,w)Y(a,z)=0,$
which is equivalent to that
(4.34) $Y(a,z)Y(a_{0}b,w)\sim\left(\frac{w-z-3\hbar}{w-z-\hbar}\right)\!Y(a_{0}b,w)Y(a,z)$
and ${\rm Sing}_{z}Y(a,z)a_{0}b=0$. With (4.31), by Lemma 4.3 we have
$x_{1}^{-1}\delta\left(\frac{w-\hbar}{x_{1}}\right)\!Y(a_{0}b,w)=Y(a,x_{1})Y(b,w)-\left(\frac{w-x_{1}+\hbar}{w-x_{1}-\hbar}\right)\!Y(b,w)Y(a,x_{1}).$
Then using commutation relations (4.30) and (4.31) we get
$(z-w+\hbar)(z-w-\hbar)Y(a,z)Y(a_{0}b,w)=(z-w+3\hbar)(z-w-\hbar)Y(a_{0}b,w)Y(a,z),$
which implies (4.34). Now, it follows. ∎
Proposition 4.10.
Let $V$ be an $\hbar$-adic nonlocal vertex algebra and let $a,b\in V$. If
(4.35) $(z-w-2\hbar)Y(a,z)Y(a,w)=(z-w+2\hbar)Y(a,w)Y(a,z),$
(4.36) $(z-w+\hbar)Y(a,z)Y(b,w)=(z-w-\hbar)Y(b,w)Y(a,z),$
and if the order-$2$ Serre relation for $(Y(a,z_{1}),Y(a,z_{2}),Y(b,w))$ holds, then for any $V$-module $(W,Y_{W})$,
(4.37) $(z-w-2\hbar)Y_{W}(a,z)Y_{W}(a,w)=(z-w+2\hbar)Y_{W}(a,w)Y_{W}(a,z),$
(4.38) $(z-w+\hbar)Y_{W}(a,z)Y_{W}(b,w)=(z-w-\hbar)Y_{W}(b,w)Y_{W}(a,z),$
and the order-$2$ Serre relation for $(Y_{W}(a,z_{1}),Y_{W}(a,z_{2}),Y_{W}(b,w))$ holds. The converse is also true for any faithful $V$-module $(W,Y_{W})$.
Proof.
In view of Proposition 4.5, under the assumptions (4.35) and (4.36), the order-$2$ Serre relation for $(Y(a,z_{1}),Y(a,z_{2}),Y(b,w))$ holds if and only if
(4.39) $Y(a,z)Y(a_{0}b,w)=Y(a_{0}b,w)Y(a,z)\left(\frac{w-z-3\hbar}{w-z-\hbar}\right)\!.$
For the same reason, under the assumption that (4.37) and (4.38) hold, the order-$2$ Serre relation for $(Y_{W}(a,z_{1}),Y_{W}(a,z_{2}),Y_{W}(b,w))$ holds if and only if
(4.40) $Y_{W}(a,z)Y_{W}(a_{0}b,w)=Y_{W}(a_{0}b,w)Y_{W}(a,z)\left(\frac{w-z-3\hbar}{w-z-\hbar}\right)\!.$
Then it follows immediately from Proposition 3.22. ∎
As an immediate consequence of Proposition 4.10 we have:
Corollary 4.11.
Let $W$ be a topologically free $\mathbb{C}$$\hbar$]]-module and let $a(x),b(x)\in U$, where $U$ is an $\hbar$-adically compatible subset of ${\mathcal{E}}_{\hbar}(W)$, such that
(4.41) $(x_{1}-x_{2}-2\hbar)a(x_{1})a(x_{2})=(x_{1}-x_{2}+2\hbar)a(x_{2})a(x_{1}),$
(4.42) $(x_{1}-x_{2}+\hbar)a(x_{1})b(x_{2})=(x_{1}-x_{2}-\hbar)b(x_{2})a(x_{1}).$
Assume that the order-$2$ Serre relation for $(a(z_{1}),a(z_{2}),b(w))$ holds. Then
$(z_{1}-z_{2}-2\hbar)Y_{{\mathcal{E}}}(a(x),z_{1})Y_{{\mathcal{E}}}(a(x),z_{2})=(z_{1}-z_{2}+2\hbar)Y_{{\mathcal{E}}}(a(x),z_{2})Y_{{\mathcal{E}}}(a(x),z_{1}),$
$(z_{1}-z_{2}+\hbar)Y_{{\mathcal{E}}}(a(x),z_{1})Y_{{\mathcal{E}}}(b(x),z_{2})=(z_{1}-z_{2}-\hbar)Y_{{\mathcal{E}}}(b(x),z_{2})Y_{{\mathcal{E}}}(a(x),z_{1})$
on ${\langle}U{\rangle}$, and the order-$2$ Serre relation for $(Y_{{\mathcal{E}}}(a(x),z_{1}),Y_{{\mathcal{E}}}(a(x),z_{2}),Y_{{\mathcal{E}}}(b(x),w))$ holds.
Recall the following standard formal series convention (see [FLM]):
(4.43) $\log(1+f(x))=\sum_{n\geq 1}(-1)^{n-1}\frac{1}{n}f(x)^{n}\in xK[[x]],$
(4.44) $e^{f(x)}=\sum_{n\geq 0}\frac{1}{n!}f(x)^{n}\in K[[x]]$
for $f(x)\in xK[[x]]$, where $K$ is any associative algebra with identity over $\mathbb{C}$}.
For $f(x,z)\in{\mathbb{C}}(x,z)$, as a convention we write
(4.45) $f(x,\hbar):=\left(\iota_{x,z}f(x,z)\right)|_{z=\hbar}\in{\mathbb{C}}((x))((\hbar)).$
Let $p(x,z),q(x,z)\in{\mathbb{C}}[x,z]$ with $p(x,0)=q(x,0)\neq 0$. Then $\frac{p(x,\hbar)}{q(x,\hbar)}=1+hf(x,\hbar)$ for some $f(x,\hbar)\in{\mathbb{C}}((x))[[\hbar]]$, and hence
(4.46) $\log\left(\frac{p(x,\hbar)}{q(x,\hbar)}\right)\in\hbar{\mathbb{C}}((x))[[\hbar]].$
In particular, for $\mu,\nu\in\hbar{\mathbb{C}}[[\hbar]],\ a\in K$, we have
(4.47) $\log\left(\frac{x+\mu}{x+\nu}\right)$ $=\log(1+\mu x^{-1})-\log(1+\nu x^{-1})=(e^{\mu\partial_{x}}-e^{\nu\partial_{x}})\log x,$
(4.48) $\left(\frac{x+\mu}{x+\nu}\right)^{a}=\exp\!\left(\!a\log\left(\frac{x+\mu}{x+\nu}\right)\!\right)\!=\exp\left(a(e^{\mu\partial_{x}}-e^{\nu\partial_{x}})\log x\right)\!.$
Next, we present a vertex-operator iterate formula. First, we establish the following technical result (cf. [JKLT2]):
Lemma 4.12.
Let $W$ be a topologically free $\mathbb{C}$$\hbar$]]-module and let
$\alpha(x)\in\operatorname{Hom}(W,W\widehat{\otimes}{\mathbb{C}}[x,x^{-1}][[\hbar]]),\ \ \beta(x)\in(\operatorname{End}W)[[x]],\ \ \gamma(x)\in{\mathbb{C}}((x))[[\hbar]],$
satisfying the condition
(4.49) $[\alpha(x_{1}),\alpha(x_{2})]=0=[\beta(x_{1}),\beta(x_{2})],\quad[\alpha(x_{1}),\beta(x_{2})]=\gamma(x_{1}-x_{2}).$
Then $U:=\{\alpha(x),\beta(x),1_{W}\}$ is an $\mathcal{S}$}-local subset of ${\mathcal{E}}_{\hbar}(W)$. Furthermore, assume $\alpha(x),\beta(x)\in\hbar(\operatorname{End}W)[[x,x^{-1}]]$ and set $E_{\gamma}=\operatorname{Res}_{z}z^{-1}\gamma(z)\in\hbar^{2}{\mathbb{C}}[[\hbar]]$. Then
(4.50) $\exp\left(\left(\alpha(x)+\beta(x)\right)_{-1}\right)1_{W}=\exp\left(\frac{1}{2}E_{\gamma}\right)\exp\alpha(x)\exp\beta(x).$
Proof.
From (4.49) it can be readily seen that $U$ is $\mathcal{S}$}-local. Then we have an $\hbar$-adic weak quantum vertex algebra ${\langle}U{\rangle}$. With (4.49), by Lemma 3.46 we get
$[Y_{\mathcal{E}}(\alpha(x),x_{1}),Y_{\mathcal{E}}(\beta(x),x_{2})]=\gamma(x_{1}-x_{2}),$
which in particular implies
$[\alpha(x)_{-1},\beta(x)_{-1}]=E_{\gamma}.$
Then by the Baker-Campbell-Hausdorff formula we have
(4.51) $\exp\left((\alpha(x)+\beta(x))_{-1}\right)=\exp\left(\frac{1}{2}E_{\gamma}\right)\exp\left(\beta(x)_{-1}\right)\exp\left(\alpha(x)_{-1}\right)\!.$
On the other hand, notice that
$\alpha(x_{1})\alpha(x_{2})^{n},\,\beta(x_{1})u(x_{2})\in{\mathcal{E}}_{\hbar}^{(2)}(W)\quad\textnormal{{for }}n\in{\mathbb{N}},\,\,u(x)\in{\langle}U{\rangle}.$
Then from the definition of $Y_{\mathcal{E}}$ we get
$\alpha(x)_{-1}\alpha(x)^{n}=\alpha(x)^{n+1}\quad\textnormal{{for }}n\in{\mathbb{N}},$
$\beta(x)_{-1}u(x)=\beta(x)u(x)\quad\textnormal{{for }}u(x)\in{\langle}U{\rangle}.$
Using (4.51) and these relations we obtain
$\exp\left((\alpha(x)+\beta(x))_{-1}\right)1_{W}$
$=$ $\exp\left({\frac{1}{2}}E_{\gamma}\right)\exp\left(\beta(x)_{-1}\right)\exp\left(\alpha(x)_{-1}\right)1_{W}$
$=$ $\exp\left({\frac{1}{2}}E_{\gamma}\right)\exp\left(\beta(x)_{-1}\right)\exp\left(\alpha(x)\right)$
$=$ $\exp\left({\frac{1}{2}}E_{\gamma}\right)\exp\left(\beta(x)\right)\exp\left(\alpha(x)\right)\!,$
as desired. ∎
Introduce a formal power series
(4.52) $L(x)=\sum_{n\geq 1}\frac{1}{n!}x^{n-1}=\frac{e^{x}-1}{x}\in{\mathbb{C}}[[x]].$
For $v\in V$ (an $\hbar$-adic nonlocal vertex algebra), denote by $Y^{+}(v,x)$ and $Y^{-}(v,x)$ the regular and singular parts of $Y(v,x)$, respectively. The following is the vertex-operator iterate formula that we need:
Proposition 4.13.
Let $V$ be an $\hbar$-adic nonlocal vertex algebra, $(W,Y_{W})$ a $V$-module. Assume $a\in V$ such that
(4.53) $[Y^{\pm}(a,x_{1}),Y^{\pm}(a,x_{2})]=0,\quad[Y^{-}(a,x_{1}),Y^{+}(a,x_{2})]=\gamma(x_{1}-x_{2}),$
where $\gamma(x)\in x^{-1}{\mathbb{C}}[x^{-1}][[\hbar]]$. Let $z\in\hbar{\mathbb{C}}[[\hbar]]$ and set
$E^{\pm}(a,z)=\exp\left(\sum_{n\in\pm{\mathbb{Z}}_{+}}\frac{1}{n}a_{n}z^{-n}\right)$
on $V$ and $W$, where $Y(a,x)=\sum_{n\in{\mathbb{Z}}}a_{n}x^{-n-1}$. Then
(4.54) $Y_{W}(E^{-}(-a,z){\bf 1},x)=\exp\left(zL(z\partial_{x})Y_{W}^{+}(a,x)\right)\exp\left(zL(z\partial_{x})Y_{W}^{-}(a,x)\right)$
$\quad=\left(1+\frac{z}{x}\right)^{a_{0}}E^{-}(-a,x+z)E^{-}(a,x)E^{+}(-a,x+z)E^{+}(a,x).$
Proof.
Set $v=\sum_{n\geq 1}\frac{1}{n}z^{n}a_{-n}{\bf 1}\in\hbar V$. We have
(4.55) $v=\sum_{n\geq 1}\frac{1}{n!}z^{n}\mathcal{D}^{n-1}a\quad\text{ and }\quad v_{-1}=\sum_{n\geq 1}\frac{1}{n}z^{n}a_{-n}.$
Then $E^{-}(-a,z)=\exp v_{-1}$. On the other hand, as
$Y_{W}(v,x)=\sum_{n\geq 1}z^{n}\frac{1}{n!}Y_{W}(\mathcal{D}^{n-1}a,x)=\sum_{n\geq 1}z^{n}\frac{1}{n!}\left(\frac{d}{dx}\right)^{n-1}Y_{W}(a,x),$
we get
(4.56) $Y_{W}^{\pm}(v,x)=\sum_{n\geq 1}z^{n}\frac{1}{n!}\left(\frac{d}{dx}\right)^{n-1}Y_{W}^{\pm}(a,x)=zL(z\partial_{x})Y_{W}^{\pm}(a,x).$
Note that the commutation relations in (4.53) amount to
$Y(a,x_{1}),Y(a,x_{2})]=\gamma(x_{1}-x_{2})-\gamma(x_{2}-x_{1}).$
In view of Lemma 3.25, we have
$Y_{W}(a,x_{1}),Y_{W}(a,x_{2})]=\gamma(x_{1}-x_{2})-\gamma(x_{2}-x_{1}),$
which amounts to
(4.57) $[Y_{W}^{\pm}(a,x_{1}),Y^{\pm}_{W}(a,x_{2})]=0,\quad[Y_{W}^{-}(a,x_{1}),Y_{W}^{+}(a,x_{2})]=\gamma(x_{1}-x_{2})$
as $\gamma(x)\in x^{-1}{\mathbb{C}}[x^{-1}][[\hbar]]$. Then using (4.56) we get
(4.58) $[Y_{W}^{-}(v,x_{1}),Y_{W}^{+}(v,x_{2})]$ $\ =z^{2}L(z\partial_{x_{1}})L(z\partial_{x_{2}})\gamma(x_{1}-x_{2})$
$\ =z^{2}L(z\partial_{x_{1}})L(-z\partial_{x_{1}})\gamma(x_{1}-x_{2})$
$\ =\tilde{\gamma}(x_{1}-x_{2}),$
where $\tilde{\gamma}(x)=z^{2}L(z\partial_{x})L(-z\partial_{x})\gamma(x)\in\hbar^{2}x^{-1}{\mathbb{C}}[x^{-1}][[\hbar]].$ Then using Lemma 4.12 with $\alpha(x)=Y_{W}^{-}(v,x)$, $\beta(x)=Y_{W}^{+}(v,x)$, and using the fact that $Y_{W}(\cdot,x)$ is a nonlocal vertex algebra homomorphism, we obtain
$Y_{W}(E^{-}(-a,z){\bf 1},x)=Y_{W}((\exp v_{-1}){\bf 1},x)=\exp\left(Y_{W}(v,x)_{-1}\right)1_{W}$
$=\exp(Y_{W}^{+}(v,x))\exp(Y_{W}^{-}(v,x))=\exp\left(zL(z\partial_{x})Y_{W}^{+}(a,x)\right)\exp\left(zL(z\partial_{x})Y_{W}^{-}(a,x)\right)\!,$
noticing that $E_{\tilde{\gamma}}:=z^{2}\operatorname{Res}_{x}x^{-1}L(z\partial_{x})L(-z\partial_{x})\gamma(x)=0$. Furthermore, we have
(4.59) $zL(z\partial_{x})Y_{W}^{+}(a,x)=(e^{z\partial_{x}}-1)\sum_{n\geq 1}\frac{1}{n}a_{-n}x^{n}=\sum_{n\geq 1}\frac{1}{n}a_{-n}\left((x+z)^{n}-x^{n}\right)\!,$
$zL(z\partial_{x})Y_{W}^{-}(a,x)$ $=(e^{z\partial_{x}}-1)\sum_{n\geq 1}(-1)\frac{1}{n}a_{n}x^{-n}+a_{0}(e^{z\partial_{x}}-1)\log x$
$=-\sum_{n\geq 1}\frac{1}{n}a_{n}\left((x+z)^{-n}-x^{-n}\right)+a_{0}\log\left(1+\frac{z}{x}\right)\!.$
Then it follows immediately. ∎
5. Double Yangians and $\hbar$-adic quantum vertex algebras
In this section, for any simply-laced generalized Cartan matrix (GCM) $A$, we introduce an algebra $\widehat{\mathcal{DY}}(A)$, generalizing centrally extended double Yangians. As the main results, we give a new field presentation of $\widehat{\mathcal{DY}}(A)$ and a (tautological) construction of a universal vacuum $\widehat{\mathcal{DY}}(A)$-module $\mathcal{V}_{A}(\ell)$ of any level $\ell\in{\mathbb{C}}$, and we prove that $\mathcal{V}_{A}(\ell)$ has a canonical $\hbar$-adic weak quantum vertex algebra structure with any (other) restricted $\widehat{\mathcal{DY}}(A)$-modules of level $\ell$ as modules.
5.1. Centrally extended double Yangian
Let $A=[a_{i,j}]_{i,j\in I}$ be a simply-laced GCM, which is fixed throughout this section. Note that $A$ is symmetric and
$a_{i,j}\in\{-1,0,2\}\quad\text{ for }i,j\in I.$
Now, we introduce the main object for this paper.
Definition 5.1.
Define $\widehat{\mathcal{DY}}(A)$ to be the topological associative unital algebra over $\mathbb{C}$$\hbar$]], generated by
(5.1) ${\left.\left\{{H_{i,n},\,X_{i,n}^{\pm}}\,\right|\,{i\in I,\,n\in{\mathbb{Z}}}\right\}}\cup\{\kappa\},$
subject to a set of relations written in terms of generating functions:
(5.2) $H_{i}^{+}(z)=1+2\hbar\sum_{n\geq 0}H_{i,n}z^{-n-1},\quad H_{i}^{-}(z)=1+2\hbar\sum_{n\geq 1}H_{i,-n}z^{n-1},$
(5.3) $\hskip 99.58464ptX_{i}^{\pm}(z)=\sum_{n\in{\mathbb{Z}}}X_{i,n}^{\pm}z^{-n-1}.$
In addition to that $\kappa$ and $H_{i,0}$ for $i\in I$ are central, the following are the defining relations with $i,j\in I$:
$\textnormal{{(DY1)}}\quad[H_{i}^{\pm}(z),H_{j}^{\pm}(w)]=0,$
$\textnormal{{(DY2)}}\quad H_{i}^{+}(z)H_{j}^{-}(w)=H_{j}^{-}(w)H_{i}^{+}(z)\frac{(z-w-a_{i,j}\hbar-\kappa\hbar)(z-w+a_{i,j}\hbar+\kappa\hbar)}{(z-w+a_{i,j}\hbar-\kappa\hbar)(z-w-a_{i,j}\hbar+\kappa\hbar)},$
$\textnormal{{(DY3)}}\quad H_{i}^{+}(z)X_{j}^{\pm}(w)=X_{j}^{\pm}(w)H_{i}^{+}(z)\left(\frac{z-w+a_{i,j}\hbar\pm{\frac{1}{2}}\kappa\hbar}{z-w-a_{i,j}\hbar\pm{\frac{1}{2}}\kappa\hbar}\right)^{\pm 1},$
$\textnormal{{(DY4)}}\quad H_{i}^{-}(z)X_{j}^{\pm}(w)=X_{j}^{\pm}(w)H_{i}^{-}(z)\left(\frac{w-z-a_{j,i}\hbar\pm{\frac{1}{2}}k\hbar}{w-z+a_{j,i}\hbar\pm{\frac{1}{2}}\kappa\hbar}\right)^{\pm 1},$
$\textnormal{{(DY5)}}\quad[X_{i}^{+}(z),X_{j}^{-}(w)]=\delta_{i,j}\frac{1}{2\hbar}\Bigg(\!H_{i}^{+}(w+{\frac{1}{2}}\kappa\hbar)z^{-1}\delta\!\left(\frac{w+\kappa\hbar}{z}\right)$
$\qquad\qquad\qquad\qquad-H_{i}^{-}(w-{\frac{1}{2}}\kappa\hbar)z^{-1}\delta\!\left(\frac{w-\kappa\hbar}{z}\right)\!\Bigg),$
$\textnormal{{(DY6)}}\quad(z-w\mp a_{i,j}\hbar)X_{i}^{\pm}(z)X_{j}^{\pm}(w)=(z-w\pm a_{i,j}\hbar)X_{j}^{\pm}(w)X_{i}^{\pm}(z),$
$\textnormal{{(DY7)}}\quad X_{i}^{\pm}(z)X_{j}^{\pm}(w)=X_{j}^{\pm}(w)X_{i}^{\pm}(z)\ \ \text{ if }a_{i,j}=0,$
$\textnormal{{(DY8)}}\quad\sum_{\sigma\in S_{2}}\Big(X_{i}^{\pm}(z_{\sigma(1)})X_{i}^{\pm}(z_{\sigma(2)})X_{j}^{\pm}(w)-2X_{i}^{\pm}(z_{\sigma(1)})X_{j}^{\pm}(w)X_{i}^{\pm}(z_{\sigma(2)})$
$\qquad\qquad+X_{j}^{\pm}(w)X_{i}^{\pm}(z_{\sigma(1)})X_{i}^{\pm}(z_{\sigma(2)})\Big)=0\quad\textnormal{{if }}a_{i,j}=-1.$
Note that relation (DY1) is equivalent to that
(5.4) $[H_{i,m},H_{j,n}]=0\quad\textnormal{{ for }}m,n\in{\mathbb{Z}}\textnormal{{ with either }}m,n\geq 0\textnormal{{ or }}m,n<0.$
Remark 5.2.
Note that if $A$ is of finite type with ${\mathfrak{g}}={\mathfrak{g}}(A)$, this definition except (DY7) is the same as in [DK, page 17] (after a parameter-change $\hbar\rightarrow 2i\hbar$). On the other hand, if ${\mathfrak{g}}=\mathfrak{sl}_{n+1}$, $\widehat{\mathcal{DY}}(A)$ is isomorphic to the algebra $\mathcal{DY}_{\hbar}({\mathfrak{g}})$ defined in [I, Corollary 3.4].
Definition 5.3.
A $\widehat{\mathcal{DY}}(A)$-module $W$ is called a restricted module if $W$ is a topologically free $\mathbb{C}$$\hbar$]]-module such that for every $i\in I,\ w\in W$,
(5.5) $\lim_{n\rightarrow\infty}H_{i,n}w=0=\lim_{n\rightarrow\infty}X_{i,n}^{\pm}w.$
Note that $H_{i}^{\pm}(z),X_{i}^{\pm}(z)\in{\mathcal{E}}_{\hbar}(W)$ $(i\in I)$ for any restricted $\widehat{\mathcal{DY}}(A)$-module $W$.
Definition 5.4.
Denote by $\widetilde{\mathcal{A}}$ the associative algebra over $\mathbb{C}$}, generated by elements
$\tilde{\kappa},\ \tilde{H}_{i,n},\ \tilde{X}_{i,n}^{\pm}\ \ (\text{where }i\in I,\ n\in{\mathbb{Z}}),$
subject to the relation (condition) that $\tilde{\kappa}$ and $\tilde{H}_{i,0}$ for $i\in I$ are central.
Consider the $\mathbb{C}$$\hbar$]]-algebra $\widetilde{\mathcal{A}}$$\hbar$]] and formulate generating functions $\tilde{H}_{i}^{\pm}(z),\ \tilde{X}_{i}^{\pm}(z)$ for $i\in I$ in the same way as for $H_{i}^{\pm}(z),\ X_{i}^{\pm}(z)$. Note that $\tilde{H}_{i}^{+}(z)$ and $\tilde{H}_{i}^{-}(z)$ are invertible elements of $\widetilde{\mathcal{A}}[[z^{-1}]][[\hbar]]\ (=\widetilde{\mathcal{A}}[[\hbar]][[z^{-1}]])$ and $\widetilde{\mathcal{A}}[[z]][[\hbar]]\ (=\widetilde{\mathcal{A}}[[\hbar]][[z]])$, respectively.
A restricted $\widetilde{\mathcal{A}}$$\hbar$]]-module is defined to be an $\widetilde{\mathcal{A}}$$\hbar$]]-module $W$ which is a topologically free $\mathbb{C}$$\hbar$]]-module such that
(5.6) $\lim_{n\rightarrow\infty}\tilde{H}_{i,n}w=0=\lim_{n\rightarrow\infty}\tilde{X}_{i,n}^{\pm}w\quad\text{ for }i\in I,\ w\in W.$
Set $q=e^{\hbar}\in{\mathbb{C}}[[\hbar]]$. Introduce the following power series in $x$ and $\hbar$:
(5.7) $G(x)=\frac{q^{x}-q^{-x}}{x}:=2\hbar\left(1+\sum_{n\geq 1}\frac{\hbar^{2n}}{(2n+1)!}x^{2n}\right)\in\hbar{\mathbb{C}}[[\hbar,x]],$
(5.8) $F(x)=\frac{1}{2\hbar}\left(1+\sum_{n\geq 1}\frac{\hbar^{2n}}{(2n+1)!}x^{2n}\right)^{-1}\in\hbar^{-1}{\mathbb{C}}[[x,\hbar]].$
Then G(x)=G(x)G(-x)=G(x), F(x)=F(x)F(-x)=F(x),
(5.9) $G(\partial_{z})=\frac{q^{\partial_{z}}-q^{-\partial_{z}}}{\partial_{z}},\ \ \ F(\partial_{z})=\frac{\partial_{z}}{q^{\partial_{z}}-q^{-\partial_{z}}},\ \ \text{and }\ F(\partial_{z})G(\partial_{z})=1.$
For $m\in{\mathbb{Z}}$, in terms of $G(x)$ we have
(5.10) $\log\left(\frac{x+m\hbar}{x-m\hbar}\right)=\left(\frac{q^{m\partial_{x}}-q^{-m\partial_{x}}}{\partial_{x}}\right)x^{-1}=G(\partial_{x})[m]_{q^{\partial_{x}}}x^{-1}.$
Furthermore, we have
(5.11) $\log\frac{(z-w-m\hbar-\kappa\hbar)(z-w+m\hbar+\kappa\hbar)}{(z-w+m\hbar-\kappa\hbar)(z-w-m\hbar+\kappa\hbar)}$
[…]
Definition 5.5.
Let $W$ be a restricted $\widetilde{\mathcal{A}}$$\hbar$]]-module. For $i\in I$, set
(5.12) $h_{i,\mathcal{Y}}^{\pm}(z)=\pm F(\partial_{z})\log\tilde{H}_{i}^{\pm}\big(z\pm\frac{1}{2}\tilde{\kappa}\hbar\big)\in\widetilde{\mathcal{A}}[[\hbar]][[z^{\mp 1}]].$
Furthermore, set
(5.13) $h_{i,\mathcal{Y}}(z)=h_{i,\mathcal{Y}}^{+}(z)+h_{i,\mathcal{Y}}^{-}(z),$
(5.14) $x_{i,\mathcal{Y}}^{+}(z)=\tilde{X}_{i}^{+}(z),$
(5.15) $x_{i,\mathcal{Y}}^{-}(z)=\tilde{X}_{i}^{-}(z-\tilde{\kappa}\hbar)\tilde{H}_{i}^{+}\big(z-\frac{1}{2}\tilde{\kappa}\hbar\big)^{-1}.$
Note that $h_{i,\mathcal{Y}}^{+}(z)$ and $h_{i,\mathcal{Y}}^{-}(z)$ coincide with the singular and regular parts of $h_{i,\mathcal{Y}}(z)$ with respect to variable $z$, respectively. For any restricted $\widetilde{\mathcal{A}}$$\hbar$]]-module $W$, we have
$h_{i,\mathcal{Y}}(z),\ x_{i,\mathcal{Y}}^{\pm}(z)\in{\mathcal{E}}_{\hbar}(W)\quad\textnormal{{for }}i\in I.$
As the first main result of this section, we have:
Theorem 5.6.
Let $W$ be a restricted $\widetilde{\mathcal{A}}$$\hbar$]]-module. For $i\in I$, set
(5.16) $C_{i}(x)$ $=\tilde{H}^{-}_{i}(x-\frac{3}{2}\tilde{\kappa}\hbar)\tilde{H}_{i}^{+}(x-\frac{1}{2}\tilde{\kappa}\hbar)^{-1}$
$=\exp\!\left(-G(\partial_{x})q^{-\tilde{\kappa}\partial_{x}}h_{i,\mathcal{Y}}^{-}(x)\right)\exp\!\left(-G(\partial_{x})q^{-\tilde{\kappa}\partial_{x}}h_{i,\mathcal{Y}}^{+}(x)\right)\!,$
an element of ${\mathcal{E}}_{\hbar}(W)$. Then $W$ is a restricted $\widehat{\mathcal{DY}}(A)$-module with
$\kappa=\tilde{\kappa},\ H_{i,n}=\tilde{H}_{i,n},\ X_{i,n}^{\pm}=\tilde{X}_{i,n}^{\pm}\quad\text{ for }i\in I$
if and only if the following relations hold for $i,j\in I$:
$[h_{i,\mathcal{Y}}(z),h_{j,\mathcal{Y}}(w)]=[a_{i,j}]_{q^{\partial_{w}}}[\kappa]_{q^{\partial_{w}}}\left((z-w+\hbar\kappa)^{-2}-(w-z+\hbar\kappa)^{-2}\right)\!,$
$[h_{i,\mathcal{Y}}(z),x_{j,\mathcal{Y}}^{\pm}(w)]=\pm x_{j,\mathcal{Y}}^{\pm}(w)[a_{i,j}]_{q^{\partial_{w}}}\left((z-w+\hbar\kappa)^{-1}+(w-z+\hbar\kappa)^{-1}\right)\!,$
$(z-w-a_{i,j}\hbar)x_{i,\mathcal{Y}}^{\pm}(z)x_{j,\mathcal{Y}}^{\pm}(w)=(z-w+a_{i,j}\hbar)x_{j,\mathcal{Y}}^{\pm}(w)x_{i,\mathcal{Y}}^{\pm}(z),$
$x_{i,\mathcal{Y}}^{\pm}(z)x_{j,\mathcal{Y}}^{\pm}(w)=x_{j,\mathcal{Y}}^{\pm}(w)x_{i,\mathcal{Y}}^{\pm}(z)\ \ \text{ if }a_{i,j}=0,$
$x_{i,\mathcal{Y}}^{+}(z)x_{j,\mathcal{Y}}^{-}(w)-\left(\frac{w-z+a_{i,j}\hbar}{w-z-a_{i,j}\hbar}\right)x_{j,\mathcal{Y}}^{-}(w)x_{i,\mathcal{Y}}^{+}(z)$
$\quad=\delta_{i,j}\frac{1}{2\hbar}\left(\!z^{-1}\delta\!\left(\frac{w}{z}\right)-C_{i}(w)z^{-1}\delta\!\left(\frac{w-2\kappa\hbar}{z}\right)\right)\!,$
and the order-$2$ Serre relations for $(x_{i,\mathcal{Y}}^{\pm}(z_{1}),x_{i,\mathcal{Y}}^{\pm}(z_{2}),x_{j,\mathcal{Y}}^{\pm}(w))$ hold if $a_{i,j}=-1$.
Proof.
Note that the relation for $h_{i,\mathcal{Y}}(z),h_{j,\mathcal{Y}}(w)]$ amounts to
$[h_{i,\mathcal{Y}}^{+}(z),h_{j,\mathcal{Y}}(w)]=[a_{i,j}]_{q^{\partial_{w}}}[\kappa]_{q^{\partial_{w}}}(z-w+\hbar\kappa)^{-2},$
$[h_{i,\mathcal{Y}}^{-}(z),h_{j,\mathcal{Y}}(w)]=-[a_{i,j}]_{q^{\partial_{w}}}[\kappa]_{q^{\partial_{w}}}(w-z+\hbar\kappa)^{-2}.$
These relations are furthermore equivalent to
(5.17) $[h_{i,\mathcal{Y}}^{\pm}(z),h_{j,\mathcal{Y}}^{\pm}(w)]=0,$
(5.18) $[h_{i,\mathcal{Y}}^{+}(z),h_{j,\mathcal{Y}}^{-}(w)]=[a_{i,j}]_{q^{\partial_{w}}}[\kappa]_{q^{\partial_{w}}}(z-w+\hbar\kappa)^{-2}$
for all $i,j\in I$ (recall that $a_{i,j}=a_{j,i}$).
On the other hand, for $i,j\in I$, by (5.11) we have
$\log\frac{(z-w-a_{i,j}\hbar-\kappa\hbar)(z-w+a_{i,j}\hbar+\kappa\hbar)}{(z-w+a_{i,j}\hbar-\kappa\hbar)(z-w-a_{i,j}\hbar+\kappa\hbar)}$
$=$ $\ -G(\partial_{z})G(\partial_{w})[a_{i,j}]_{q^{\partial_{z}}}[\kappa]_{q^{\partial_{z}}}(z-w)^{-2}.$
Then (DY2) amounts to
$[\log H_{i}^{+}(z),\log H_{j}^{-}(w)]=-G(\partial_{z})G(\partial_{w})[a_{i,j}]_{q^{\partial_{z}}}[\kappa]_{q^{\partial_{z}}}(z-w)^{-2},$
which is equivalent to
$[\log H_{i}^{+}(z+1/2\kappa\hbar),\log H_{j}^{-}(w-1/2\kappa\hbar)]$
[…]
This is furthermore equivalent to
$-[F(\partial_{z})\log H_{i}^{+}(z+1/2\kappa\hbar),F(\partial_{w})\log H_{j}^{-}(w-1/2\kappa\hbar)]$
$=$ $\ [a_{i,j}]_{q^{\partial_{w}}}[\kappa]_{q^{\partial_{w}}}(z-w+\kappa\hbar)^{-2}.$
Thus, (5.18) is equivalent to (DY2). It is clear that (5.17) is equivalent to (DY1).
Recall
$x_{i,\mathcal{Y}}^{+}(x)=X_{i}^{+}(x),\quad x_{i,\mathcal{Y}}^{-}(x)=X_{i}^{-}(x-\kappa\hbar)H_{i}^{+}(x-\frac{1}{2}\kappa\hbar)^{-1}.$
Rewrite (DY2) as
$H_{i}^{-}(z)H_{j}^{+}(w)^{-1}=H_{j}^{+}(w)^{-1}H_{i}^{-}(z)\frac{(w-z-a_{j,i}\hbar-\kappa\hbar)(w-z+a_{j,i}\hbar+\kappa\hbar)}{(w-z+a_{j,i}\hbar-\kappa\hbar)(w-z-a_{j,i}\hbar+\kappa\hbar)},$
which is equivalent to
$[\log H_{i}^{-}(z),H_{j}^{+}(w)^{-1}]$
[…]
Then
$[h_{i,\mathcal{Y}}^{-}(z),H_{j}^{+}(w-\frac{1}{2}\kappa\hbar)^{-1}]$
[…]
Thus
(5.19) $[h_{i,\mathcal{Y}}(z),H_{j}^{+}(w-\frac{1}{2}\kappa\hbar)^{-1}]$
[…]
Note that relation (DY3) is equivalent to
(5.20) $[\log H_{i}^{+}(z),X_{j}^{\pm}(w)]=\pm X_{j}^{\pm}(w)\log\left(\frac{z-w+a_{i,j}\hbar\pm\frac{1}{2}\kappa\hbar}{z-w-a_{i,j}\hbar\pm\frac{1}{2}\kappa\hbar}\right)\!.$
By (5.10) this amounts to
$[h_{i,\mathcal{Y}}^{+}(z),X_{j}^{+}(w)]=\$$X_{j}^{+}(w)F(\partial_{z})\log\left(\frac{z-w+a_{i,j}\hbar+\kappa\hbar}{z-w-a_{i,j}+\kappa\hbar}\right)$
[…]
$[h_{i,\mathcal{Y}}^{+}(z),X_{j}^{-}(w)]=\$$-X_{j}^{-}(w)F(\partial_{z})\log\left(\frac{z-w+a_{i,j}\hbar}{z-w-a_{i,j}\hbar}\right)$
[…]
Similarly, relation (DY4) is equivalent to
(5.21) $[\log H_{i}^{-}(z),X_{j}^{\pm}(w)]=\pm X_{j}^{\pm}(w)\log\left(\frac{w-z-a_{i,j}\hbar\pm\frac{1}{2}\kappa\hbar}{w-z+a_{i,j}\hbar\pm\frac{1}{2}\kappa\hbar}\right)\!.$
This is equivalent to
$[h_{i,\mathcal{Y}}^{-}(z),X_{j}^{+}(w)]=\$$-X_{j}^{+}(w)F(\partial_{z})\log\left(\frac{w-z-a_{i,j}\hbar+\kappa\hbar}{w-z+a_{i,j}\hbar+\kappa\hbar}\right)$
[…]
and
$[h_{i,\mathcal{Y}}^{-}(z),X_{j}^{-}(w)]=\$$X_{j}^{-}(w)F(\partial_{z})\log\left(\frac{w-z-a_{i,j}\hbar}{w-z+a_{i,j}\hbar}\right)$
[…]
Therefore, (DY3) together with (DY4) is equivalent to
(5.22) $[h_{i,\mathcal{Y}}(z),x_{j,\mathcal{Y}}^{+}(w)]$ $\ =[h_{i,\mathcal{Y}}^{+}(z)+h_{i,\mathcal{Y}}^{-}(z),X_{j}^{+}(w)]$
$\ =X_{j}^{+}(w)[a_{i,j}]_{q^{\partial_{w}}}((z-w+\hbar\kappa)^{-1}+(w-z+\hbar\kappa)^{-1})$
$\ =x_{j,\mathcal{Y}}^{+}(w)[a_{i,j}]_{q^{\partial_{w}}}((z-w+\hbar\kappa)^{-1}+(w-z+\hbar\kappa)^{-1}),$
(5.23) $[h_{i,\mathcal{Y}}(z),X_{j}^{-}(w)]=-X_{j}^{-}(w)[a_{i,j}]_{q^{\partial_{w}}}((z-w)^{-1}+(w-z)^{-1}).$
Note that the very last relation amounts to
(5.24) $[h_{i,\mathcal{Y}}(z),X_{j}^{-}(w-\kappa\hbar)]$
[…]
which in the presence of (5.19) is equivalent to
(5.25) $[h_{i,\mathcal{Y}}(z),x_{j,\mathcal{Y}}^{-}(w)]=-x_{j,\mathcal{Y}}^{-}(w)[a_{i,j}]_{q^{\partial_{w}}}((z-w+\kappa\hbar)^{-1}+(w-z+\kappa\hbar)^{-1}).$
For $r\in I$, setting
$\bar{X}_{r}(z)=X_{r}^{-}(z-\kappa\hbar),\ \ K_{r}(z)=H_{r}^{+}(z-\kappa\hbar/2)^{-1},$
we have $x_{r,\mathcal{Y}}^{-}(z)=\bar{X}_{r}(z)K_{r}(z).$ It is clear that (DY6) is equivalent to
(5.26) $(z-w+a_{r,s}\hbar)\bar{X}_{r}(z)\bar{X}_{s}(w)=(z-w-a_{r,s}\hbar)\bar{X}_{s}(w)\bar{X}_{r}(z)$
for $r,s\in I$. On the other hand, in the presence of (DY1) and (DY3) we have
(5.27) $K_{r}(z)K_{s}(w)=K_{s}(w)K_{r}(z),$
(5.28) $K_{r}(z)\bar{X}_{s}(w)=\left(\frac{z-w+a_{r,s}\hbar}{z-w-a_{r,s}\hbar}\right)\bar{X}_{s}(w)K_{r}(z).$
Furthermore, we have
(5.29) $(z-w-a_{r,s}\hbar)x_{r,\mathcal{Y}}^{-}(z)x_{s,\mathcal{Y}}^{-}(w)=(z-w+a_{r,s}\hbar)\bar{X}_{r}(z)\bar{X}_{s}(w)K_{r}(z)K_{s}(w),$
(5.30) $(z-w+a_{r,s}\hbar)x_{s,\mathcal{Y}}^{-}(w)x_{r,\mathcal{Y}}^{-}(z)=(z-w-a_{r,s}\hbar)\bar{X}_{s}(w)\bar{X}_{r}(z)K_{s}(w)K_{r}(z),$
noticing that $a_{r,s}=a_{s,r}$. Then it follows that (DY6) is equivalent to
(5.31) $(z-w-a_{r,s}\hbar)x_{r,\mathcal{Y}}^{-}(z)x_{s,\mathcal{Y}}^{-}(w)=(z-w+a_{r,s}\hbar)x_{s,\mathcal{Y}}^{-}(w)x_{r,\mathcal{Y}}^{-}(z).$
Now, we discuss (DY7). It is clear that (DY7) holds if and only if
(5.32) $\bar{X}_{i}(z)\bar{X}_{j}(w)=\bar{X}_{j}(w)\bar{X}_{i}(z)\quad\mbox{if }a_{i,j}=0.$
As $K_{r}(z)\in\mathcal{A}[[\hbar]][[z^{-1}]]$, (5.28) is equivalent to
(5.33) $\left(\frac{z-w-a_{r,s}\hbar}{z-w+a_{r,s}\hbar}\right)K_{r}(z)\bar{X}_{s}(w)=\bar{X}_{s}(w)K_{r}(z).$
Then using (5.28) we have
$x_{j,\mathcal{Y}}^{-}(w)x_{i,\mathcal{Y}}^{-}(z)$ $=\bar{X}_{j}(w)\bar{X}_{i}(z)K_{j}(w)K_{i}(z),$
$x_{i,\mathcal{Y}}^{-}(z)x_{j,\mathcal{Y}}^{-}(w)$ $=\bar{X}_{i}(z)K_{i}(z)\bar{X}_{j}(w)K_{j}(w)$
$=\bar{X}_{i}(z)\bar{X}_{j}(w)K_{i}(z)K_{j}(w),\quad\mbox{since }a_{i,j}=0.$
Thus, in the presence of (DY1) and (DY3), (5.32) is equivalent to
$x_{j,\mathcal{Y}}^{-}(w)x_{i,\mathcal{Y}}^{-}(z)=x_{i,\mathcal{Y}}^{-}(z)x_{j,\mathcal{Y}}^{-}(w)\quad\mbox{if }a_{i,j}=0.$
This gives the equivalence involving (DY7).
Next, we consider the Serre relations. As $x_{r,\mathcal{Y}}^{+}(z)=X_{r}^{+}(z)$ for $r\in I$, we only need to consider the Serre relation for $(x_{i,\mathcal{Y}}^{-}(z_{1}),x_{i,\mathcal{Y}}^{-}(z_{2}),x_{j,\mathcal{Y}}^{-}(w))$. With relation (5.26), by Lemma 4.6 we have
$(z-w)^{-1}(z-w+2\hbar)\bar{X}_{i}(z)\bar{X}_{i}(w)=(w-z)^{-1}(w-z+2\hbar)\bar{X}_{i}(w)\bar{X}_{i}(z).$
It follows that
${\mathopen{\overset{\bullet}{{{\bullet}}}}{\bar{X}_{i}(z_{1})\bar{X}_{i}(z_{2})\bar{X}_{j}(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}={\mathopen{\overset{\bullet}{{{\bullet}}}}{\bar{X}_{i}(z_{2})\bar{X}_{i}(z_{1})\bar{X}_{j}(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}.$
Then using (5.31) and (5.29) we have
$z_{1}^{-1}\delta\left(\frac{w+\hbar}{z_{1}}\right)z_{2}^{-1}\delta\left(\frac{w-\hbar}{z_{2}}\right){\mathopen{\overset{\bullet}{{{\bullet}}}}{x_{i,\mathcal{Y}}^{-}(z_{1})x_{i,\mathcal{Y}}^{-}(z_{2})x_{j,\mathcal{Y}}^{-}(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}$
[…]
Applying $\operatorname{Res}_{z_{1}}\operatorname{Res}_{z_{2}}$ we get
${\mathopen{\overset{\bullet}{{{\bullet}}}}{x_{i,\mathcal{Y}}^{-}(z_{1})x_{i,\mathcal{Y}}^{-}(z_{2})x_{j,\mathcal{Y}}^{-}(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}|_{z_{1}=w+\hbar,z_{2}=w-\hbar}$
$=$ ${\mathopen{\overset{\bullet}{{{\bullet}}}}{X_{i}^{-}(z_{1})X_{i}^{-}(z_{2})X_{j}^{-}(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}|_{z_{1}=w-\hbar,z_{2}=w+\hbar}K_{i}(w-\hbar)K_{i}(w+\hbar)K_{j}(w).$
From this we see that ${\mathopen{\overset{\bullet}{{{\bullet}}}}{x_{i,\mathcal{Y}}^{-}(z_{1})x_{i,\mathcal{Y}}^{-}(z_{2})x_{j,\mathcal{Y}}^{-}(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}|_{z_{1}=w+\hbar,z_{2}=w-\hbar}=0$ if and only if
${\mathopen{\overset{\bullet}{{{\bullet}}}}{X_{i}^{-}(z_{1})X_{i}^{-}(z_{2})X_{j}^{-}(w)}\mathclose{\overset{\bullet}{{{\bullet}}}}}|_{z_{1}=w-\hbar,z_{2}=w+\hbar}=0.$
Then by Proposition 4.8, the Serre relation for $(\bar{X}_{i}(z_{1}),\bar{X}_{i}(z_{2}),\bar{X}_{j}(w))$ holds if and only if the Serre relation for $(x_{i,\mathcal{Y}}^{-}(z_{1}),x_{i,\mathcal{Y}}^{-}(z_{2}),x_{j,\mathcal{Y}}^{-}(w))$ holds.
Finally, we consider relation (DY5). A variation of this is
$[X_{i}^{+}(z),\bar{X}_{j}(w)]=\delta_{i,j}\frac{1}{2\hbar}\left(\!K_{i}(w)^{-1}z^{-1}\delta\!\left(\frac{w}{z}\right)-H_{i}^{-}(w-3/2\kappa\hbar)z^{-1}\delta\!\left(\frac{w-2\kappa\hbar}{z}\right)\right)\!.$
From (DY3) we have
$X_{i}^{+}(z)K_{j}(w)=K_{j}(w)X_{i}^{+}(z)\left(\frac{w-z+a_{j,i}\hbar}{w-z-a_{j,i}\hbar}\right)\!.$
Then
$x_{i,\mathcal{Y}}^{+}(z)x_{j,\mathcal{Y}}^{-}(w)-\left(\frac{w-z+a_{j,i}\hbar}{w-z-a_{j,i}\hbar}\right)x_{j,\mathcal{Y}}^{-}(w)x_{i,\mathcal{Y}}^{+}(z)$
[…]
$=\$$[X_{i}^{+}(z),\bar{X}_{j}(w)]K_{j}(w).$
On the other hand, we have
$\delta_{i,j}\frac{1}{2\hbar}\left(\!z^{-1}\delta\!\left(\frac{w}{z}\right)-C_{i}(w)z^{-1}\delta\!\left(\frac{w-2\kappa\hbar}{z}\right)\!\right)\!$
[…]
This proves the equivalence regarding (DY5). Now, the proof is complete. ∎
With Theorem 5.6, from now on we shall freely use $x_{i,\mathcal{Y}}^{\pm}(z),\ h_{i,\mathcal{Y}}(z)$ $(i\in I)$ as generating functions for $\widehat{\mathcal{DY}}(A)$. For $i\in I$, write
(5.34) $x_{i,\mathcal{Y}}^{\pm}(z)=\sum_{n\in{\mathbb{Z}}}x_{i,\mathcal{Y}}^{\pm}(n)z^{-n-1},\quad h_{i,\mathcal{Y}}(z)=\sum_{n\in{\mathbb{Z}}}h_{i,\mathcal{Y}}(n)z^{-n-1}.$
Definition 5.7.
Let $\ell\in{\mathbb{C}}$. A vacuum $\widehat{\mathcal{DY}}(A)$-module of level $\ell$ is a pair $(W,w_{0})$, where $W$ is a restricted $\widehat{\mathcal{DY}}(A)$-module of level $\ell$ and $w_{0}$ is a vector in $W$ such that
(5.35) $W=(\widehat{\mathcal{DY}}(A)w_{0})[[\hbar]]^{\prime},$
(5.36) $x_{i,\mathcal{Y}}^{\pm}(n)w_{0}=0=h_{i,\mathcal{Y}}(n)w_{0}\quad\text{ for }i\in I,\ n\geq 0.$
Let $W$ be any restricted $\widehat{\mathcal{DY}}(A)$-module. Identify $\kappa$ with its corresponding element of $\operatorname{End}W$ and set
(5.37) $U_{W}={\left.\left\{{h_{i,\mathcal{Y}}(z),x_{i,\mathcal{Y}}^{\pm}(z)}\,\right|\,{i\in I}\right\}}\cup\{1_{W},\kappa\}.$
It follows from Theorem 5.6 and Lemmas 3.46 and 3.47 that $U_{W}$ is an $\mathcal{S}$-local subset of ${\mathcal{E}}_{\hbar}(W)$, which is $\hbar$-adically compatible by Lemma 3.40. Then we have an $\hbar$-adic nonlocal vertex algebra ${\langle}U_{W}{\rangle}$ with $W$ as a faithful module. Furthermore, we have:
Proposition 5.8.
Let $W$ be any restricted $\widehat{\mathcal{DY}}(A)$-module. Then the $\hbar$-adic nonlocal vertex algebra ${\langle}U_{W}{\rangle}$ is also a restricted $\widehat{\mathcal{DY}}(A)$-module with $\kappa$ acting as its associated left multiplication on ${\langle}U_{W}{\rangle}$ and with
$h_{i,\mathcal{Y}}(z_{0})=Y_{\mathcal{E}}(h_{i,\mathcal{Y}}(z),z_{0}),\quad x_{i,\mathcal{Y}}^{\pm}(z_{0})=Y_{\mathcal{E}}(x_{i,\mathcal{Y}}^{\pm}(z),z_{0})\quad\text{ for }i\in I.$
Furthermore, if $W$ is of level $\ell\in{\mathbb{C}}$, then ${\langle}U_{W}{\rangle}$ is a $\widehat{\mathcal{DY}}(A)$-module of level $\ell$.
Proof.
First, identify $\kappa$ with its associated left multiplication, an operator on ${\mathcal{E}}_{\hbar}(W)$. Then $Y_{{\mathcal{E}}}(\kappa,z)=\kappa$. As $\kappa$ is a central element of $\widehat{\mathcal{DY}}(A)$, it follows that
$[\kappa,Y_{{\mathcal{E}}}(h_{i,\mathcal{Y}}(z),z_{2})]=0=[\kappa,Y_{{\mathcal{E}}}(x_{i,\mathcal{Y}}^{\pm}(z),z_{2})]\quad\text{ for }i\in I.$
With the relations in Theorem 5.6, by Lemmas 3.26 and 3.25 we have
$[Y_{{\mathcal{E}}}(h_{i,\mathcal{Y}}(z),z_{1}),Y_{{\mathcal{E}}}(h_{j,\mathcal{Y}}(z),z_{2})]$
$=\$$[a_{i,j}]_{q^{\frac{\partial}{\partial{z_{2}}}}}[\kappa]_{q^{\frac{\partial}{\partial{z_{2}}}}}\left((z_{1}-z_{2}+\kappa\hbar)^{-2}-(z_{2}-z_{1}+\kappa\hbar)^{-2}\right)\!,$
$[Y_{{\mathcal{E}}}(h_{i,\mathcal{Y}}(z),z_{1}),Y_{{\mathcal{E}}}(x_{j,\mathcal{Y}}^{\pm}(z),z_{2}]$
$=\$$\pm Y_{{\mathcal{E}}}(x_{j,\mathcal{Y}}^{\pm}(z),z_{2})[a_{i,j}]_{q^{\frac{\partial}{\partial{z_{2}}}}}\left((z_{1}-z_{2}+\kappa\hbar)^{-1}-(-z_{2}+z_{1}-\kappa\hbar)^{-1}\right)\!,$
$(z_{1}-z_{2}-a_{i,j}\hbar)Y_{{\mathcal{E}}}(x_{i,\mathcal{Y}}^{\pm}(z),z_{1})Y_{{\mathcal{E}}}(x_{j,\mathcal{Y}}^{\pm}(z),z_{2})$
$=\$$(z_{1}-z_{2}+a_{i,j}\hbar)Y_{{\mathcal{E}}}(x_{j,\mathcal{Y}}^{\pm}(z),z_{2})Y_{{\mathcal{E}}}(x_{i,\mathcal{Y}}^{\pm}(z),z_{1})$
for $i,j\in I$ and
$Y_{{\mathcal{E}}}(x_{i,\mathcal{Y}}^{\pm}(z),z_{1})Y_{{\mathcal{E}}}(x_{j,\mathcal{Y}}^{\pm}(z),z_{2})=Y_{{\mathcal{E}}}(x_{j,\mathcal{Y}}^{\pm}(z),z_{2})Y_{{\mathcal{E}}}(x_{i,\mathcal{Y}}^{\pm}(z),z_{1})$
if $a_{i,j}=0$. In case $a_{i,j}=-1$, by Corollary 4.11 the order-$2$ Serre relations hold for the triples $(Y_{{\mathcal{E}}}(x_{i,\mathcal{Y}}^{\pm}(z),z_{1}),Y_{{\mathcal{E}}}(x_{i,\mathcal{Y}}^{\pm}(z),z_{2}),Y_{{\mathcal{E}}}(x_{j,\mathcal{Y}}^{\pm}(z),w))$.
Now, we consider the remaining commutator relation. For $i,j\in I$, if $i\neq j$, by Lemma 3.47 we have
$Y_{{\mathcal{E}}}(x_{i,\mathcal{Y}}^{+}(z),z_{1})Y_{{\mathcal{E}}}(x_{j,\mathcal{Y}}^{-}(z),z_{2})-\!\left(\frac{z_{2}-z_{1}+a_{i,j}\hbar}{z_{2}-z_{1}-a_{i,j}\hbar}\right)\!Y_{{\mathcal{E}}}(x_{j,\mathcal{Y}}^{-}(z),z_{2})Y_{{\mathcal{E}}}(x_{i,\mathcal{Y}}^{+}(z),z_{1})=0.$
Assume $i=j$. Recall
$C_{i}(x)=\exp\!\left(-G(\partial_{x})q^{-\kappa\partial_{x}}h_{i,\mathcal{Y}}^{-}(x)\right)\exp\!\left(-G(\partial_{x})q^{-\kappa\partial_{x}}h_{i,\mathcal{Y}}^{+}(x)\right)\!.$
Notice that
$z^{-1}\delta\!\left(\frac{w-2\kappa\hbar}{z}\right)\!C_{i}(w)=\sum_{n\geq 0}(-2\kappa\hbar)^{n}C_{i}(w)\frac{1}{n!}\left(\frac{\partial}{\partial w}\right)^{n}z^{-1}\delta\!\left(\frac{w}{z}\right)\!.$
By Proposition 3.22 we have
$C_{i}(z)=1_{W}-2\hbar x_{i,\mathcal{Y}}^{+}(z)_{0}x_{i,\mathcal{Y}}^{-}(z)\in{\langle}U_{W}{\rangle}$
and
(5.38) $Y_{{\mathcal{E}}}(x_{i,\mathcal{Y}}^{+}(z),z_{1})Y_{{\mathcal{E}}}(x_{i,\mathcal{Y}}^{-}(z),z_{2})-\!\left(\frac{z_{2}-z_{1}+2\hbar}{z_{2}-z_{1}-2\hbar}\right)\!Y_{{\mathcal{E}}}(x_{i,\mathcal{Y}}^{-}(z),z_{2})Y_{{\mathcal{E}}}(x_{i,\mathcal{Y}}^{+}(z),z_{1})$
$=\$$\frac{1}{2\hbar}\!\left(\!z_{1}^{-1}\delta\!\left(\frac{z_{2}}{z_{1}}\right)-Y_{{\mathcal{E}}}\left(C_{i}(z),z_{2}\right)z_{1}^{-1}\delta\!\left(\frac{z_{2}-2\kappa\hbar}{z_{1}}\right)\!\right)\!.$
Recall that $h_{i,\mathcal{Y}}(x)\in{\langle}U_{W}{\rangle}$. In the following, we use Proposition 4.13 to show
(5.39) $C_{i}(x)=q^{(1-\kappa)\mathcal{D}}E^{-}(-h_{i,\mathcal{Y}}(x),-2\hbar)1_{W},$
where $\mathcal{D}$ denotes the $D$-operator of ${\langle}U_{W}{\rangle}$.
Note that
$Y_{{\mathcal{E}}}^{\pm}(h_{i,\mathcal{Y}}(x),z),Y_{{\mathcal{E}}}^{\pm}(h_{i,\mathcal{Y}}(x),w)]=0,$
$Y_{{\mathcal{E}}}^{-}(h_{i,\mathcal{Y}}(x),z),Y_{{\mathcal{E}}}^{+}(h_{i,\mathcal{Y}}(x),w)]=[2]_{q^{\partial_{w}}}[\kappa]_{q^{\partial_{w}}}(z-w+\kappa\hbar)^{-2}.$
With $W$ naturally a ${\langle}U_{W}{\rangle}$-module, by Proposition 4.13 we have
(5.40) $Y_{W}(E^{-}(-h_{i,\mathcal{Y}}(z),\zeta){\bf 1},x)$
$=$ $\exp\left(\zeta L(\zeta\partial_{x})Y_{W}^{+}(h_{i,\mathcal{Y}}(z),x)\right)\exp(\zeta L(\zeta\partial_{x})Y_{W}^{-}(h_{i,\mathcal{Y}}(z),x))$
for $\zeta\in\hbar{\mathbb{C}}[[\hbar]]$, where $L(x)=\frac{e^{x}-1}{x}\in{\mathbb{C}}[[x]]$. Note that
(5.41) $-G(x)q^{-x}=\frac{e^{-2\hbar x}-1}{x}=-2\hbar L(-2\hbar x).$
With $Y_{W}^{\pm}(h_{i,\mathcal{Y}}(z),x)=h_{i,\mathcal{Y}}^{\mp}(x)$, using (5.40) with $\zeta=-2\hbar$ we get
$Y_{W}\!\left(q^{(1-\kappa)\mathcal{D}}E^{-}(-h_{i,\mathcal{Y}}(z),-2\hbar)1_{W},x\right)\!$
[…]
Then (5.39) follows as $W$ is a faithful module.
Now, using (5.39) and using Proposition 4.13 for the adjoint ${\langle}U_{W}{\rangle}$-module we get
$Y_{{\mathcal{E}}}(C_{i}(z),x)$
[…]
Finally, by invoking Theorem 5.6 we conclude that ${\langle}U_{W}{\rangle}$ is a restricted $\widehat{\mathcal{DY}}(A)$-module with $h_{i,\mathcal{Y}}(z_{0})=Y_{\mathcal{E}}(h_{i,\mathcal{Y}}(z),z_{0})$, $x_{i,\mathcal{Y}}^{\pm}(z_{0})=Y_{\mathcal{E}}(x_{i,\mathcal{Y}}^{\pm}(z),z_{0})$ for $i\in I$.
The second assertion is clear. ∎
5.2. Universal vacuum $\widehat{\mathcal{DY}}(A)$-module ${\mathcal{V}}_{A}(\ell)$
Let $\ell\in{\mathbb{C}}$. We here construct a universal vacuum $\widehat{\mathcal{DY}}(A)$-module ${\mathcal{V}}_{A}(\ell)$ of level $\ell$ and present the second main result of this paper.
First, let $\widetilde{\mathcal{U}}$ denote the associative algebra with identity over $\mathbb{C}$}, generated by set
$\{\tilde{\kappa}\}\cup\{\tilde{e}_{i}(m),\tilde{f}_{i}(m),\tilde{h}_{i}(m)\ |\ i\in I,\ m\in{\mathbb{Z}}\}.$
For $X\in\{\tilde{e}_{i},\tilde{f}_{i},\tilde{h}_{i}\ |\ i\in I\}$, form a generating function
(5.42) $X(z)=\sum_{m\in{\mathbb{Z}}}X(m)z^{-m-1}.$
Second, consider the $\mathbb{C}$$\hbar$]]-algebra $\widetilde{\mathcal{U}}$$\hbar$]], and define $\widetilde{J}_{A,\ell}$ to be the two-sided ideal, generated by element $\tilde{\kappa}-\ell$ and by the following relations for $i,j\in I$:
(5.43) $\tilde{h}_{i}(z)\tilde{h}_{j}(w)-\tilde{h}_{j}(w)\tilde{h}_{i}(z)=[a_{i,j}]_{q^{\partial_{w}}}[\ell]_{q^{\partial_{w}}}((z-w+\ell\hbar)^{-2}-(w-z+\ell\hbar)^{-2}),$
(5.44) $\tilde{h}_{i}(z)\tilde{e}_{j}(w)-\tilde{e}_{j}(w)\tilde{h}_{i}(z)=\tilde{e}_{j}(w)[a_{i,j}]_{q^{\partial_{w}}}((z-w+\ell\hbar)^{-1}-(w-z+\ell\hbar)^{-1}),$
(5.45) $\tilde{h}_{i}(z)\tilde{f}_{j}(w)-\tilde{f}_{j}(w)\tilde{h}_{i}(z)=-\tilde{f}_{j}(w)[a_{i,j}]_{q^{\partial_{w}}}((z-w+\ell\hbar)^{-1}-(w-z+\ell\hbar)^{-1}),$
(5.46) $(z-w-a_{i,j}\hbar)\tilde{e}_{i}(z)\tilde{e}_{j}(w)=(z-w+a_{i,j}\hbar)\tilde{e}_{j}(w)\tilde{e}_{i}(z),$
(5.47) $(z-w-a_{i,j}\hbar)\tilde{f}_{i}(z)\tilde{f}_{j}(w)=(z-w+a_{i,j}\hbar)\tilde{f}_{j}(w)\tilde{f}_{i}(z),$
(5.48) $(z-w+a_{i,j}\hbar)\tilde{e}_{i}(z)\tilde{f}_{j}(w)=(z-w-a_{i,j}\hbar)\tilde{f}_{j}(w)\tilde{e}_{i}(z)\quad\text{ if }i\neq j,$
(5.49) $(z-w)(z-w+2\ell\hbar)\!\left((z-w+2\hbar)\tilde{e}_{i}(z)\tilde{f}_{i}(w)-(z-w-2\hbar)\tilde{f}_{i}(w)\tilde{e}_{i}(z)\right)\!=0,$
and the order-$2$ Serre relations for $(\tilde{e}_{i}(z_{1}),\tilde{e}_{i}(z_{1}),\tilde{e}_{j}(w))$ and $(\tilde{f}_{i}(z_{1}),\tilde{f}_{i}(z_{1}),\tilde{f}_{j}(w))$ for $i,j\in I$ with $a_{i,j}=-1$.
Note that for any $r,s\in{\mathbb{Z}}$, the coefficients of monomial $z^{r}w^{s}$ in all the relations above indeed lie in $\widetilde{\mathcal{U}}$$\hbar$]].
In view of Lemma 3.7, $\overline{[\widetilde{J}_{A,\ell}]}$ is a two-sided ideal of $\widetilde{\mathcal{U}}$$\hbar$]] and a topologically free $\mathbb{C}$$\hbar$]]-submodule. Then set
(5.50) $\mathcal{U}_{A,\ell}=\widetilde{\mathcal{U}}[[\hbar]]/\overline{[\widetilde{J}_{A,\ell}]},$
which is an associative $\mathbb{C}$$\hbar$]]-algebra and a topologically free $\mathbb{C}$$\hbar$]]-module by Lemma 3.8. For $i\in I,\ m\in{\mathbb{Z}}$, set
(5.51) $\bar{e}_{i}(m)=\tilde{e}_{i}(m)+\overline{[\widetilde{J}_{A,\ell}]},\ \bar{f}_{i}(m)=\tilde{f}_{i}(m)+\overline{[\widetilde{J}_{A,\ell}]},\ \bar{h}_{i}(m)=\tilde{h}_{i}(m)+\overline{[\widetilde{J}_{A,\ell}]}\in\mathcal{U}_{A,\ell}.$
Then define generating functions $\bar{e}_{i}(z),\bar{f}_{i}(z),\bar{h}_{i}(z)$ for $i\in I$ correspondingly.
Remark 5.9.
Note that from Theorem 5.6, every restricted $\widehat{\mathcal{DY}}(A)$-module of level $\ell$ is naturally a restricted $\mathcal{U}_{A,\ell}$-module.
Third, let $\mathcal{J}_{A,\ell}$ be the left ideal of $\mathcal{U}_{A,\ell}$, generated by the subset
$\{\bar{e}_{i}(n),\bar{f}_{i}(n),\bar{h}_{i}(n)\ |\ i\in I,\ n\geq 0\}.$
Then $\overline{[\mathcal{J}_{A,\ell}]}$ is a topologically free $\mathbb{C}$$\hbar$]]-submodule and a left ideal of $\mathcal{U}_{A,\ell}$. Define
(5.52) $\mathcal{V}_{A,\ell}=\mathcal{U}_{A,\ell}/\overline{[\mathcal{J}_{A,\ell}]},$
which is a (left) $\mathcal{U}_{A,\ell}$-module and a topologically free $\mathbb{C}$$\hbar$]]-module. Set
${\bf 1}=1+\overline{[\mathcal{J}_{A,\ell}]}\in\mathcal{V}_{A,\ell},$
and furthermore set
(5.53) $\hat{e}_{i}=\bar{e}_{i}(-1){\bf 1},\ \ \hat{f}_{i}=\bar{f}_{i}(-1){\bf 1},\ \ \hat{h}_{i}=\bar{h}_{i}(-1){\bf 1}\in\mathcal{V}_{A,\ell}\quad\text{for }i\in I.$
As a key result of this section, we have:
Proposition 5.10.
There exists an $\hbar$-adic weak quantum vertex algebra structure on $\mathcal{V}_{A,\ell}$, which is uniquely determined by the condition that $\bf 1$} is the vacuum vector and
(5.54) $Y(\hat{e}_{i},z)=\bar{e}_{i}(z),\ \ Y(\hat{f}_{i},z)=\bar{f}_{i}(z),\ \ Y(\hat{h}_{i},z)=\bar{h}_{i}(z)\quad\text{for }i\in I.$
Proof.
First of all, it is clear that $\mathcal{V}_{A,\ell}$ is topologically spanned over $\mathbb{C}$$\hbar$]] by vectors $X_{1}\cdots X_{r}{\bf 1}$ for $r\geq 0,\ X_{j}\in\{\bar{e}_{i}(m),\bar{f}_{i}(m),\bar{h}_{i}(m)\ |\ i\in I,\ m\in{\mathbb{Z}}\}.$ From the construction of $\mathcal{V}_{A,\ell}$ as a $\mathcal{U}_{A,\ell}$-module, we have
$\bar{e}_{i}(x){\bf 1},\ \bar{f}_{i}(x){\bf 1},\ \bar{h}_{i}(x){\bf 1}\in\mathcal{V}_{A,\ell}[[x]]\subset(\mathcal{V}_{A,\ell})_{\hbar}((x))\ \ \text{ for }i\in I.$
Here, we need to show that $\mathcal{V}_{A,\ell}$ is a restricted $\mathcal{U}_{A,\ell}$-module. Indeed, it follows from induction and technical Lemmas 5.19 and 5.20 that for every $w\in\mathcal{V}_{A,\ell}$,
$\bar{e}_{i}(x)w,\ \bar{f}_{i}(x)w,\ \bar{h}_{i}(x)w\in(\mathcal{V}_{A,\ell})_{\hbar}((x))\ \ \text{ for }i\in I.$
That is,
(5.55) $\bar{e}_{i}(x),\ \bar{f}_{i}(x),\ \bar{h}_{i}(x)\in{\mathcal{E}}_{\hbar}(\mathcal{V}_{A,\ell})\ \ \text{ for }i\in I.$
Furthermore, by Lemma 3.45 $U:=\{\bar{e}_{i}(x),\bar{f}_{i}(x),\bar{h}_{i}(x)\ |\ i\in I\}$ is an $\mathcal{S}$}-local subset of ${\mathcal{E}}_{\hbar}(\mathcal{V}_{A,\ell})$. Then $U$ generates an $\hbar$-adic nonlocal vertex algebra ${\langle}U{\rangle}$ inside ${\mathcal{E}}_{\hbar}(\mathcal{V}_{A,\ell})$ with $\mathcal{V}_{A,\ell}$ as a faithful module. It follows from Lemmas 3.46 and 3.47 that ${\langle}U{\rangle}$ is a $\mathcal{U}_{A,\ell}$-module with $\bar{e}_{i}(m),\bar{f}_{i}(m),\bar{h}_{i}(m)$ for $i\in I,\ m\in{\mathbb{Z}}$ acting as $\bar{e}_{i}(x)_{m},\bar{f}_{i}(x)_{m},\bar{h}_{i}(x)_{m}$, respectively, and it is clear that ${\langle}U{\rangle}$ is a restricted $\mathcal{U}_{A,\ell}$-module. Note that
$u(x)_{n}1_{\mathcal{V}_{A,\ell}}=0\ \ \text{ for all }u(x)\in{\langle}U{\rangle},\ n\in{\mathbb{N}}.$
It follows from the construction of $\mathcal{V}_{A,\ell}$ and Lemma 3.6 that there is a $\mathcal{U}_{A,\ell}$-module morphism $\phi:\mathcal{V}_{A,\ell}\rightarrow{\langle}U{\rangle}$ such that $\phi({\bf 1})=1_{\mathcal{V}_{A,\ell}}$. Now, by invoking Theorem 3.41 we conclude that there exists an $\hbar$-adic weak quantum vertex algebra structure on $\mathcal{V}_{A,\ell}$ with the required condition. ∎
Furthermore, using Theorem 3.42 we immediately have:
Proposition 5.11.
For any restricted $\mathcal{U}_{A,\ell}$-module $W$, there exists a $\mathcal{V}_{A,\ell}$-module structure $Y_{W}(\cdot,z)$ which is uniquely determined by
(5.56) $Y_{W}(\hat{e}_{i},z)=\bar{e}_{i}(z),\ \ Y_{W}(\hat{f}_{i},z)=\bar{f}_{i}(z),\ \ Y_{W}(\hat{h}_{i},z)=\bar{h}_{i}(z)\ \ \text{ for }i\in I.$
On the other hand, for any $\mathcal{V}_{A,\ell}$-module $(W,Y_{W})$, there exists a restricted $\mathcal{U}_{A,\ell}$-module structure on $W$ such that
(5.57) $\bar{e}_{i}(z)=Y_{W}(\hat{e}_{i},z),\ \ \bar{f}_{i}(z)=Y_{W}(\hat{f}_{i},z),\ \ \bar{h}_{i}(z)=Y_{W}(\hat{h}_{i},z)\ \ \text{ for }i\in I.$
We continue to define a quotient of the $\hbar$-adic weak quantum vertex algebra $\mathcal{V}_{A,\ell}$, which will be a vacuum $\widehat{\mathcal{DY}}(A)$-module of level $\ell$. For $i\in I$, set
(5.58) $C_{i}=e^{(1-\ell)\hbar\mathcal{D}}E^{-}(-\hat{h}_{i},-2\hbar){\bf 1}\in\mathcal{V}_{A,\ell},$
where $\mathcal{D}$ is the canonical $D$-operator of $\mathcal{V}_{A,\ell}$ and
$E^{-}(-\hat{h}_{i},-2\hbar)=\exp\!\left(\sum_{n\in{\mathbb{Z}}_{+}}\frac{1}{n}\hat{h}_{i}(-n)(-2\hbar)^{n}\right)\!.$
Note that $C_{i}={\bf 1}-2\hbar\hat{h}_{i}+O(\hbar^{2})$ (as $\mathcal{D}{\bf 1}=0$).
Denote by $K$ the subset of $\mathcal{V}_{A,\ell}$, consisting of vectors
(5.59) $(\hat{e}_{i})_{0}(\hat{f}_{j})-\frac{1}{2\hbar}\delta_{i,j}({\bf 1}-C_{i}),\ \ (\hat{e}_{i})_{n}(\hat{f}_{j})-\delta_{i,j}\ell(-2\ell\hbar)^{n-1}C_{i}\ \ \text{ for }i,j\in I,\ n\geq 1$
and vectors
(5.60) $(\hat{e}_{i})_{n}(\hat{e}_{j}),\ (\hat{f}_{i})_{n}(\hat{f}_{j})\quad\text{ for }i,j\in I\ \text{with }a_{i,j}=0\ \text{and for }n\geq 0.$
Definition 5.12.
Let $(K)$ denote the strong ideal of $\mathcal{V}_{A,\ell}$, generated by $K$. Set
(5.61) ${\mathcal{V}}_{A}(\ell)=\mathcal{V}_{A,\ell}/(K),$
which is a topologically free $\mathbb{C}$$\hbar$]]-module and an $\hbar$-adic weak quantum vertex algebra.
For $i\in I$, set
(5.62) $\check{e}_{i}=\hat{e}_{i}+(K),\ \check{f}_{i}=\hat{f}_{i}+(K),\ \check{h}_{i}=\hat{h}_{i}+(K)\in{\mathcal{V}}_{A}(\ell),$
and furthermore set
(5.63) $\mathfrak{c}_{i}=e^{(1-\ell)\hbar\mathcal{D}}E^{-}(-\check{h}_{i},-2\hbar){\bf 1}\in{\mathcal{V}}_{A}(\ell),$
where $\mathcal{D}$ is the $D$-operator of ${\mathcal{V}}_{A}(\ell)$. Then we have:
Theorem 5.13.
Let $\ell\in{\mathbb{C}}$. Then there exists a restricted $\widehat{\mathcal{DY}}(A)$-module structure of level $\ell$ on $\mathcal{V}_{A}(\ell)$, which is uniquely determined by
(5.64) $x_{i,\mathcal{Y}}^{+}(z)=Y(\check{e}_{i},z),\ \ x_{i,\mathcal{Y}}^{-}(z)=Y(\check{f}_{i},z),\ \ h_{i,\mathcal{Y}}(z)=Y(\check{h}_{i},z)\ \ \text{ for }i\in I.$
Furthermore, $(\mathcal{V}_{A}(\ell),{\bf 1})$ is a universal vacuum $\widehat{\mathcal{DY}}(A)$-module of level $\ell$.
Proof.
We first prove the existence. By Lemma 3.44, relations (5.48) and (5.49) imply
$Y(\check{e}_{i},z)Y(\check{f}_{j},w)\sim\left(\frac{w-z+a_{i,j}\hbar}{w-z-a_{i,j}\hbar}\right)\!Y(\check{f}_{j},w)Y(\check{e}_{i},z).$
Then the $\mathcal{S}$}-commutator formula gives
(5.65) $Y(\check{e}_{i},z)Y(\check{f}_{j},w)-\!\left(\frac{w-z+a_{i,j}\hbar}{w-z-a_{i,j}\hbar}\right)\!Y(\check{f}_{j},w)Y(\check{e}_{i},z)$
[…]
From the construction of ${\mathcal{V}}_{A}(\ell)$, we have
(5.66) $(\check{e}_{i})_{0}(\check{f}_{j})=\frac{1}{2\hbar}\delta_{i,j}({\bf 1}-\mathfrak{c}_{i}),\ \ (\check{e}_{i})_{n}(\check{f}_{j})=\delta_{i,j}\ell(-2\ell\hbar)^{n-1}\mathfrak{c}_{i}$
for $i,j\in I,\ n\geq 1$. Combining this with (5.65) we get
(5.67) $Y(\check{e}_{i},z)Y(\check{f}_{j},w)-\!\left(\frac{w-z+a_{i,j}\hbar}{w-z-a_{i,j}\hbar}\right)\!Y(\check{f}_{j},w)Y(\check{e}_{i},z)$
[…]
Note that
$G(x)q^{-\ell x}=2\hbar L(-2\hbar x)q^{(1-\ell)x}.$
Using Proposition 4.13 we obtain
(5.68) $Y(\mathfrak{c}_{i},x)$
[…]
Let $i,j\in I$ with $a_{i,j}=0$. We have $(\check{e}_{i})_{n}(\check{e}_{j})=0=(\check{f}_{i})_{n}(\check{f}_{j})$ for $n\geq 0$. With relations (5.46) and (5.47), using the arguments above we get
(5.69) $Y(\check{e}_{i},z)Y(\check{e}_{j},w)=Y(\check{e}_{j},w)Y(\check{e}_{i},z),$
(5.70) $Y(\check{f}_{i},z)Y(\check{f}_{j},w)=Y(\check{f}_{j},w)Y(\check{f}_{i},z).$
Then it follows from Theorem 5.6 that ${\mathcal{V}}_{A}(\ell)$ is a $\widehat{\mathcal{DY}}(A)$-module of level $\ell$ with
$x_{i,\mathcal{Y}}^{+}(z)=Y(\check{e}_{i},z),\ \ x_{i,\mathcal{Y}}^{-}(z)=Y(\check{f}_{i},z),\ \ h_{i,\mathcal{Y}}(z)=Y(\check{h}_{i},z)\ \ \text{ for }i\in I.$
The uniqueness and vacuum module assertions are clear.
Next, we prove that $({\mathcal{V}}_{A}(\ell),{\bf 1})$ is universal. Let $W$ be any restricted $\widehat{\mathcal{DY}}(A)$-module of level $\ell$ with a vacuum vector $w_{0}$. Set
$U_{W}=\{x_{i,\mathcal{Y}}^{+}(z),\ x_{i,\mathcal{Y}}^{-}(z),\ h_{i,\mathcal{Y}}(z)\ |\ i\in I\}\cup\{1_{W}\}.$
Then $U_{W}$ is an $\mathcal{S}$}-local subset of ${\mathcal{E}}_{\hbar}(W)$, which generates an $\hbar$-adic nonlocal vertex algebra ${\langle}U_{W}{\rangle}$ with $W$ as a module. By Lemma 3.24, there is a ${\langle}U_{W}{\rangle}$-module morphism $\psi_{0}:{\langle}U_{W}{\rangle}\rightarrow W$ with $\psi_{0}(1_{W})=w_{0}$. Note that ${\langle}U_{W}{\rangle}$ is a restricted $\widehat{\mathcal{DY}}(A)$-module of level $\ell$ by Proposition 5.8. It is easy to see that $\psi_{0}$ is also a $\widehat{\mathcal{DY}}(A)$-module morphism.
On the other hand, as ${\langle}U_{W}{\rangle}$ is a $\mathcal{U}_{A,\ell}$-module by Remark 5.9, there exists a $\mathcal{U}_{A,\ell}$-module morphism $\hat{\psi}_{z}:\mathcal{U}_{A,\ell}\rightarrow{\langle}U_{W}{\rangle}$ such that $\hat{\psi}_{z}(1)=1_{W}$. With ${\langle}U_{W}{\rangle}$ torsion free and separated, by Lemma 3.6 $\hat{\psi}_{z}$ reduces to a $\mathcal{U}_{A,\ell}$-module morphism $\tilde{\psi}_{z}:\mathcal{V}_{A,\ell}\rightarrow{\langle}U_{W}{\rangle}$ with $\tilde{\psi}_{z}({\bf 1})=1_{W}$. Since $\mathcal{V}_{A,\ell}$ as a $\mathcal{U}_{A,\ell}$-module is generated from $\bf 1$}, it follows that $\tilde{\psi}_{z}$ is a morphism of $\hbar$-adic nonlocal vertex algebras. Hence, $W$ is a $\mathcal{V}_{A,\ell}$-module via $\tilde{\psi}_{z}$. By reversing the arguments in the existence part above, we see that $K\subset\ker Y_{W}(\cdot,z)=\ker(\tilde{\psi}_{z})$. Then $\tilde{\psi}_{z}$ reduces to an $\hbar$-adic nonlocal vertex algebra morphism $\psi_{1}:{\mathcal{V}}_{A}(\ell)\rightarrow{\langle}U_{W}{\rangle}$. Consequently, $\psi:=\psi_{0}\psi_{1}$ is a $\widehat{\mathcal{DY}}(A)$-module morphism from ${\mathcal{V}}_{A}(\ell)$ to $W$ with $\psi({\bf 1})=w_{0}$. ∎
In summary, for any $\ell\in{\mathbb{C}}$, ${\mathcal{V}}_{A}(\ell)$ is an $\hbar$-adic weak quantum vertex algebra and a restricted $\widehat{\mathcal{DY}}(A)$-module of level $\ell$ with
(5.71) $Y(\check{e}_{i},z)=x_{i,\mathcal{Y}}^{+}(z),\ \ Y(\check{f}_{i},z)=x_{i,\mathcal{Y}}^{-}(z),\ \ Y(\check{h}_{i},z)=h_{i,\mathcal{Y}}(z)\ \ \text{ for }i\in I.$
With the notations
$Y(u,z)=\sum_{m\in{\mathbb{Z}}}u_{m}z^{-m-1}\quad\text{for }u\in\{\check{e}_{i},\ \check{f}_{i},\ \check{h}_{i}\ |\ i\in I\},$
$X(z)=\sum_{m\in{\mathbb{Z}}}X(m)z^{-m-1}\quad\text{for }X\in\{x_{i,\mathcal{Y}}^{\pm},\ h_{i,\mathcal{Y}}\ |\ i\in I\},$
we have
(5.72) $(\check{e}_{i})_{m}=x_{i,\mathcal{Y}}^{+}(m),\ \ (\check{f}_{i})_{m}=x_{i,\mathcal{Y}}^{-}(m),\ \ (\check{h}_{i})_{m}=h_{i,\mathcal{Y}}(m)$
for $i\in I,\ m\in{\mathbb{Z}}.$ For convenience, locally set $\check{e}_{i}^{+}=\check{e}_{i}$ and $\check{e}_{i}^{-}=\check{f}_{i}$.
Remark 5.14.
The following relations hold on ${\mathcal{V}}_{A}(\ell)$ for $i,j\in I$:
$[Y(\check{h}_{i},z),Y(\check{h}_{j},w)]=[a_{i,j}]_{q^{\partial_{w}}}[\ell]_{q^{\partial_{w}}}\left((z-w+\ell\hbar)^{-2}-(w-z+\ell\hbar)^{-2}\right)\!,$
$[Y(\check{h}_{i},z),Y(\check{e}_{j}^{\pm},w)]=\pm Y(\check{e}_{j}^{\pm},w)[a_{i,j}]_{q^{\partial_{w}}}\left((z-w+\ell\hbar)^{-1}+(w-z+\ell\hbar)^{-1}\right)\!,$
$(z-w-a_{i,j}\hbar)Y(\check{e}_{i}^{\pm},z)Y(\check{e}_{j}^{\pm},w)=(z-w+a_{i,j}\hbar)Y(\check{e}_{j}^{\pm},w)Y(\check{e}_{i}^{\pm},z),$
$Y(\check{e}_{i}^{\pm},z)Y(\check{e}_{j}^{\pm},w)=Y(\check{e}_{j}^{\pm},w)Y(\check{e}_{i}^{\pm},z)\quad\text{ if }a_{i,j}=0,$
$Y(\check{e}_{i},z)Y(\check{f}_{j},w)-\left(\frac{w-z+a_{i,j}\hbar}{w-z-a_{i,j}\hbar}\right)\!Y(\check{f}_{j},w)Y(\check{e}_{i},z)$
$\quad=\delta_{i,j}\frac{1}{2\hbar}\!\left(\!z^{-1}\delta\!\left(\frac{w}{z}\right)-Y\!\left(q^{(1-\ell)\mathcal{D}}E^{-}(-\check{h}_{i},-2\hbar){\bf 1},w\right)z^{-1}\delta\!\left(\frac{w-2\ell\hbar}{z}\right)\!\right)\!,$
where $\mathcal{D}$ denotes the $D$-operator of ${\mathcal{V}}_{A}(\ell)$. In addition, for $i,j\in I$ with $a_{i,j}=-1$, the Serre relations for $(Y(\check{e}_{i}^{\pm},z_{1}),Y(\check{e}_{i}^{\pm},z_{2}),Y(\check{e}_{j}^{\pm},w))$ hold. On the other hand, for any ${\mathcal{V}}_{A}(\ell)$-module $(W,Y_{W})$, all of these relations hold on $W$ by Propositions 3.26 and 3.22 and by Lemma 3.25. Furthermore, we have
(5.73) $Y_{W}\!\left(e^{(1-\ell)\hbar\mathcal{D}}E^{-}(-\check{h}_{i},-2\hbar){\bf 1},x\right)$
[…]
Theorem 5.15.
Let $\ell\in{\mathbb{C}}$. For any restricted $\widehat{\mathcal{DY}}(A)$-module $W$ of level $\ell$, there exists a ${\mathcal{V}}_{A}(\ell)$-module structure $Y_{W}(\cdot,z)$ on $W$, which is uniquely determined by
(5.74) $Y_{W}(\check{e}_{i},z)=x_{i,\mathcal{Y}}^{+}(z),\ \ Y_{W}(\check{f}_{i},z)=x_{i,\mathcal{Y}}^{-}(z),\ \ Y_{W}(\check{h}_{i},z)=h_{i,\mathcal{Y}}(z)\ \ \text{ for }i\in I.$
On the other hand, for any ${\mathcal{V}}_{A}(\ell)$-module $(W,Y_{W})$, there exists a restricted $\widehat{\mathcal{DY}}(A)$-module structure of level $\ell$ on $W$ such that
(5.75) $x_{i,\mathcal{Y}}^{+}(z)=Y_{W}(\check{e}_{i},z),\ \ x_{i,\mathcal{Y}}^{-}(z)=Y_{W}(\check{f}_{i},z),\ \ h_{i,\mathcal{Y}}(z)=Y_{W}(\check{h}_{i},z)\ \ \text{ for }i\in I.$
Proof.
For the first assertion, set
$U=\{\check{e}_{i},\ \check{f}_{i},\ \check{h}_{i}\ |\ i\in I\}\cup\{{\bf 1}\}\subset{\mathcal{V}}_{A}(\ell),$
and define a map $Y_{W}^{0}(\cdot,z):U\rightarrow{\mathcal{E}}_{\hbar}(W)$ by $Y_{W}^{0}({\bf 1},z)=1_{W}$,
$Y_{W}^{0}(\check{e}_{i},z)=x_{i,\mathcal{Y}}^{+}(z),\ \ Y_{W}^{0}(\check{f}_{i},z)=x_{i,\mathcal{Y}}^{-}(z),\ \ Y_{W}^{0}(\check{h}_{i},z)=h_{i,\mathcal{Y}}(z)\ \ \text{ for }i\in I.$
By Proposition 5.8, ${\langle}U_{W}{\rangle}$ is a restricted $\widehat{\mathcal{DY}}(A)$-module of level $\ell$, containing $1_{W}$ as a vacuum vector. Then by Theorem 5.13 there exists a $\widehat{\mathcal{DY}}(A)$-module homomorphism $\psi_{z}:{\mathcal{V}}_{A}(\ell)\rightarrow\langle U_{W}\rangle$ with $\psi_{z}({\bf 1})=1_{W}$. For $i\in I,\ m\in{\mathbb{Z}},\ v\in{\mathcal{V}}_{A}(\ell)$, we have
$\psi_{z}((\check{e}_{i}^{\pm})_{m}v)=\psi_{z}(x_{i,\mathcal{Y}}^{\pm}(m)v)=x_{i,\mathcal{Y}}^{\pm}(m)\psi_{z}(v)=x_{i,\mathcal{Y}}^{\pm}(z)_{m}\psi_{z}(v),$
$\psi_{z}((\check{h}_{i})_{m}v)=h_{i,\mathcal{Y}}(z)_{m}\psi_{z}(v).$
Then the first assertion follows from Theorem 3.42. The second assertion follows from Remark 5.14 and Theorem 5.6. ∎
Definition 5.16.
Define $\widehat{\mathcal{L}}_{A}$ to be the Lie algebra generated by set
(5.76) ${\left.\left\{{h_{i,m},\,e_{i,m}^{\pm}}\,\right|\,{i\in I,\ m\in{\mathbb{Z}}}\right\}}\cup\{{\bf k}\}$
with $\bf k$}, subject to the following relations for $i,j\in I$:
(L1) $[h_{i}(z),h_{j}(w)]=a_{i,j}\frac{\partial}{\partial{w}}z^{-1}\delta\!\left(\frac{w}{z}\right){\bf k},$
(L2) $[h_{i}(z),e_{j}^{\pm}(w)]=\pm a_{i,j}e_{j}^{\pm}(w)z^{-1}\delta\!\left(\frac{w}{z}\right)\!,$
(L3) $[e_{i}^{+}(z),e_{j}^{-}(w)]=\delta_{i,j}\left(\!h_{i}(w)z^{-1}\delta\!\left(\frac{w}{z}\right)+{\bf k}\frac{\partial}{\partial{w}}z^{-1}\delta\!\left(\frac{w}{z}\right)\right)\!,$
(L4) $[e_{i}^{\pm}(z),e_{j}^{\pm}(w)]=0\quad\textnormal{{if }}a_{i,j}\geq 0,$
(L5) $(z-w)[e_{i}^{\pm}(z),e_{j}^{\pm}(w)]=0\quad\textnormal{{if }}a_{i,j}=-1,$
(L6) $[e_{i}^{\pm}(z_{1}),[e_{i}^{\pm}(z_{2}),e_{j}^{\pm}(w)]]=0\quad\textnormal{{ if }}a_{i,j}=-1,$
where
$h_{i}(z)=\sum_{m\in{\mathbb{Z}}}h_{i,m}z^{-m-1},\quad e_{i}^{\pm}(z)=\sum_{m\in{\mathbb{Z}}}e_{i,m}^{\pm}z^{-m-1}.$
Lemma 5.17.
The quotient space ${\mathcal{V}}_{A}(\ell)/\hbar{\mathcal{V}}_{A}(\ell)$ is an $\widehat{\mathcal{L}}_{A}$-module of level $\ell$ with
$e_{i}^{\pm}(z)=Y(\check{e}_{i}^{\pm},z),\ \ h_{i}(z)=Y(\check{h}_{i},z)\ \ \text{ for }i\in I.$
Furthermore, ${\mathcal{V}}_{A}(\ell)/\hbar{\mathcal{V}}_{A}(\ell)$ as an $\widehat{\mathcal{L}}_{A}$-module is generated by vector ${\bf 1}+\hbar{\mathcal{V}}_{A}(\ell)$.
Proof.
For $i\in I$, as
$q^{(1-\ell)\mathcal{D}}E^{-}(-\check{h}_{i},-2\hbar){\bf 1}={\bf 1}-2\hbar\check{h}_{i}+O(\hbar^{2}),$
$z^{-1}\delta\!\left(\frac{w-2\ell\hbar}{z}\right)\!=z^{-1}\delta\!\left(\frac{w}{z}\right)-2\ell\hbar\frac{\partial}{\partial w}z^{-1}\delta\!\left(\frac{w}{z}\right)+O(\hbar^{2}),$
we have
$z^{-1}\delta\!\left(\frac{w-2\ell\hbar}{z}\right)\!q^{(1-\ell)\mathcal{D}}E^{-}(-\check{h}_{i},-2\hbar){\bf 1}$
[…]
This implies that relation (L3) holds on ${\mathcal{V}}_{A}(\ell)/\hbar{\mathcal{V}}_{A}(\ell)$. Note that in the presence of relations (L4) and (L5), (L6) is equivalent to the order-$2$ Serre relation. Then it follows from Remark 5.14 that ${\mathcal{V}}_{A}(\ell)/\hbar{\mathcal{V}}_{A}(\ell)$ is an $\widehat{\mathcal{L}}_{A}$-module of level $\ell$. It is clear that ${\mathcal{V}}_{A}(\ell)/\hbar{\mathcal{V}}_{A}(\ell)$ is generated by vector ${\bf 1}+\hbar{\mathcal{V}}_{A}(\ell)$. ∎
Assume that $A$ is of finite type. Set ${\mathfrak{g}}={\mathfrak{g}}(A)$. Then $\widehat{\mathcal{L}}_{A}$ is isomorphic to $\widehat{{\mathfrak{g}}}$ (see [G]). For $\ell\in{\mathbb{C}}$, let $V_{\widehat{{\mathfrak{g}}}}(\ell,0)$ be the universal vacuum $\widehat{{\mathfrak{g}}}-$module of level $\ell$, which is naturally a vertex algebra. It follows that there exists a surjective $\widehat{{\mathfrak{g}}}-$module homomorphism
(5.77) $\phi:\ V_{\widehat{{\mathfrak{g}}}}(\ell,0)\rightarrow{\mathcal{V}}_{A}(\ell)/\hbar{\mathcal{V}}_{A}(\ell)\ \ \text{ with }\phi({\bf 1})={\bf 1}+\hbar\mathcal{V}_{A}(\ell).$
Furthermore, $\phi$ is a vertex algebra homomorphism. From [Li3] (cf. [EK]), $V_{\widehat{{\mathfrak{g}}}}(\ell,0)$ is non-degenerate. On the other hand, if $\ell$ is generic, $V_{\widehat{{\mathfrak{g}}}}(\ell,0)$ is a simple vertex algebra. With all of these facts, we immediately have:
Proposition 5.18.
Let $A$ be of finite type and let $\ell\in{\mathbb{C}}$. Assume that the $\widehat{{\mathfrak{g}}}-$module homomorphism $\phi$ defined above is also injective, which is the case if $\ell$ is generic and if ${\mathcal{V}}_{A}(\ell)\neq 0$. Then ${\mathcal{V}}_{A}(\ell)$ is a non-degenerate $\hbar$-adic quantum vertex algebra with ${\mathcal{V}}_{A}(\ell)/\hbar{\mathcal{V}}_{A}(\ell)\cong V_{\widehat{{\mathfrak{g}}}}(\ell,0)$ as a vertex algebra.
Appendix, Two technical results
Let $W=W_{0}[[\hbar]]$ be a topologically free $\mathbb{C}$$\hbar$]]-module. Note that $(\operatorname{End}W)[[x,x^{-1}]]$ is naturally a ${\mathbb{C}}[x,x^{-1}][[\hbar]]$-module. Recall that
$W_{\hbar}((x))=W_{0}((x))[[\hbar]].$
Lemma 5.19.
Let $a(x),b(x),c(x)\in(\operatorname{End}W)[[x,x^{-1}]],\ \Psi_{1}(z),\Psi_{2}(z)\in{\mathbb{C}}[[z]],\ k,k^{\prime}\in{\mathbb{Z}}_{+}$, such that
(5.78) $[a(x_{1}),b(x_{2})]=c(x_{2})\left(\Psi_{1}(\hbar\partial_{x_{2}})(x_{1}-x_{2})^{-k}+\Psi_{2}(\hbar\partial_{x_{2}})(x_{2}-x_{1})^{-k^{\prime}}\right)\!.$
Assume $w\in W$ such that $a(x)w,b(x)w,c(x)w\in W_{\hbar}((x))$. Then $a(x)b_{m}w\in W_{\hbar}((x))$ for every $m\in{\mathbb{Z}}$, where $b(x)=\sum_{m\in{\mathbb{Z}}}b_{m}x^{-m-1}$.
Proof.
Let $m\in{\mathbb{Z}}$. We have
$a(x_{1})b_{m}w=\$$b_{m}a(x_{1})w+\operatorname{Res}_{x_{2}}x_{2}^{m}c(x_{2})w\left(\Psi_{2}(\hbar\partial_{x_{2}})(x_{2}-x_{1})^{-k^{\prime}}\right)$
$+\operatorname{Res}_{x_{2}}x_{2}^{m}c(x_{2})w\left(\Psi_{1}(\hbar\partial_{x_{2}})(x_{1}-x_{2})^{-k}\right)\!.$
Notice that the first two terms on the right-hand side lie in $W_{\hbar}((x_{1}))$. Now, consider the third term. Let $n\in{\mathbb{Z}}_{+}$. We have
$\Psi_{1}(\hbar\partial_{x_{2}})(x_{1}-x_{2})^{-k}\equiv\sum_{i=0}^{n-1}\mu_{i}(x_{1}-x_{2})^{-k-i}\hbar^{i}\ \mod\hbar^{n}{\mathbb{C}}[x_{1}^{-1}][[x_{2},\hbar]],$
where $\mu_{i}\in{\mathbb{C}}$. As $c(x)w\in W_{\hbar}((x))$, there exists $r\in{\mathbb{N}}$ such that $c_{m+j}w\in\hbar^{n}W$ for all $j>r$. Then we get
$\operatorname{Res}_{x_{2}}x_{2}^{m}c(x_{2})w\left(\Psi_{1}(\hbar\partial_{x_{2}})(x_{1}-x_{2})^{-k}\right)$
[…]
This proves $\operatorname{Res}_{x_{2}}x_{2}^{m}c(x_{2})w\left(\Psi_{1}(\hbar\partial_{x_{2}})(x_{1}-x_{2})^{-k}\right)\in W_{\hbar}((x_{1}))$. Therefore, we have $a(x)b_{m}w\in W_{\hbar}((x))$. ∎
Lemma 5.20.
Let
$a(x),b(x)\in(\operatorname{End}W)[[x,x^{-1}]],\ p(x,y),q(x,y)\in{\mathbb{C}}[x,y]$
such that $p(x,0)=x^{l}$ with $l\in{\mathbb{N}}$ and
(5.79) $p(x_{1}-x_{2},\hbar)a(x_{1})b(x_{2})=q(x_{2}-x_{1},\hbar)b(x_{2})a(x_{1}).$
Let $w\in W$ such that $a(x)w,b(x)w\in W_{\hbar}((x))$. Then $a(x)b_{m}w\in W_{\hbar}((x))$ for all $m\in{\mathbb{Z}}$.
Proof.
We must prove that for any $m\in{\mathbb{Z}}$ and for any $n\in{\mathbb{Z}}_{+}$, $\pi_{n}(a(x)b_{m}w)\in(W/\hbar^{n}W)((x))$. Let $n$ be arbitrarily fixed. From assumption, we have
$p(x_{1}-x_{2},\hbar)\pi_{n}(a(x_{1})b(x_{2})w)=q(x_{2}-x_{1},\hbar)\pi_{n}(b(x_{2})a(x_{1})w),$
$\pi_{n}(a(x)w),\pi_{n}(b(x)w)\in(W/\hbar^{n}W)((x)).$
Notice that
$p(x_{1}-x_{2},\hbar)\pi_{n}(a(x_{1})b(x_{2})w)\in(W/\hbar^{n}W)[[x_{1},x_{1}^{-1}]]((x_{2})),$
$q(x_{2}-x_{1},\hbar)\pi_{n}(b(x_{2})a(x_{1})w)\in(W/\hbar^{n}W)[[x_{2},x_{2}^{-1}]]((x_{1})).$
Consequently, we get
(5.80) $p(x_{1}-x_{2},\hbar)\pi_{n}(a(x_{1})b(x_{2})w)\in(W/\hbar^{n}W)((x_{1},x_{2})),$
while
$\pi_{n}(a(x_{1})b(x_{2})w)\in(W/\hbar^{n}W)[[x_{1},x_{1}^{-1}]]((x_{2})).$
As $p(x,0)=x^{l}$ is an invertible element of ${\mathbb{C}}[x,x^{-1}]$, $p(x,\hbar)$ is an invertible element of ${\mathbb{C}}[x,x^{-1}][[\hbar]]$, so that $p(x_{1}-x_{2},\hbar)^{-1}$ exists in ${\mathbb{C}}[x_{1},x_{1}^{-1}][[x_{2},\hbar]]$. Then multiplying (5.80) by $p(x_{1}-x_{2},\hbar)^{-1}$ we obtain
$\pi_{n}(a(x_{1})b(x_{2})w)\in(W/\hbar^{n}W)((x_{1}))((x_{2})).$
This proves that $\pi_{n}(a(x)b_{m}w)\in(W/\hbar^{n}W)((x))$ for every $m\in{\mathbb{Z}}$, as desired. ∎